### v2

# Turning Sand to Gold: Recycling Data to Bridge On-Policy and Off-Policy Learning via Causal Bound

Abstract
Deep reinforcement learning (DRL) agents excel in solving complex decision-making tasks across various domains. However, they often require a substantial number of training steps and a vast experience replay buffer, leading to significant computational and resource demands. To address these challenges, we introduce a novel theoretical result that leverages the Neyman-Rubin potential outcomes framework into DRL. Unlike most methods that focus on bounding the counterfactual loss, we establish a causal bound on the factual loss, which is analogous to the on-policy loss in DRL. This bound is computed by storing past value network outputs in the experience replay buffer, effectively utilizing data that is usually discarded. Extensive experiments across the Atari 2600 and MuJoCo domains on various agents, such as DQN and SAC, achieve up to 383% higher reward ratio, outperforming the same agents without our proposed term, and reducing the experience replay buffer size by up to 96%, significantly improving sample efficiency at a negligible cost [^1] Our code is available at https://shaham-lab.github.io/TurningSandToGold/.
1 Introduction
Deep reinforcement learning agents have demonstrated remarkable success in solving complex decision-making tasks across various areas of interest, such as gaming [1, 2, 3, 4, 5], robotics [6, 7, 8], healthcare [9, 10, 11], and autonomous driving [12, 13, 14]. It also has a vital role in enhancing large language models [15, 16]. Despite these achievements, DRL is still challenging due to its high computational and resource demands. For example, AlphaGo Zero was trained for 29 million self-play games [17], and the training experience for OpenAI’s Rubik’s Cube robotic hand is roughly equated to 13 thousand years [18].
Figure 1: Log-scaled reward improvements comparison between agents using the additional SUFT OPE term and the baseline agents without it across 57 Atari games. The results demonstrate the superior performance of our method across the majority of games. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Left: Double DQN SUFT outperforms the baseline agent in 35 out of the 40 valid games; Right: PPO SUFT outperforms the baseline agent in 39 out of the 42 valid games.
In DRL, we can categorize agents into two types: on-policy agents, like A3C and PPO [19, 20], and off-policy agents, such as DQN and SAC [21, 22]. On-policy agents are generally simpler and often considered first. They interact with the environment, collect experiences using a behavior policy, and once they optimize the target policy, discard all previously collected experiences to ensure that the behavior and target policies remain identical. This consistency leads to more stable convergence but at the cost of sample inefficiency, as each optimization iteration requires collecting new data. Off-policy agents, on the other hand, allow the agent to learn from previously collected experiences, meaning it can have different behavior and target policies. This methodology improves sample efficiency by leveraging past experiences stored in a replay buffer. While this improves sample efficiency, it can also lead to reduced learning stability due to the mismatch between the behavior and target policies. As a result, improving the learning stability typically requires a substantial number of training steps and a vast experience replay buffer, leading to significant computational and resource demands.
To mitigate these challenges, researchers have proposed numerous strategies designed to reduce computational demands and improve learning stability. A widely adopted approach focuses on enhancing the experience replay buffer sampling, aiming to improve convergence rates by optimizing the selection of experiences [23, 24, 25, 26]. While this approach improves convergence, it still involves high computational resources and does not fully address the divergence between on-policy and off-policy methods. Other works have attempted to bridge the divergence by suggesting additional computable terms, e.g., [27]. Although such approaches have improved learning stability, they often require extensive data collection and involve complex techniques.
This work addresses these limitations by introducing a novel theoretical result in causal inference and incorporating it into the DRL framework. Specifically, we establish a causal upper bound on the unobserved factual loss using the observed counterfactual loss and the estimated treatment effect. Within the DRL framework, we incorporate this as an upper bound for the on-policy loss using the standard off-policy loss and an off-policy evaluation (OPE) term, which quantifies the estimated value discrepancy between the behavior policy and the target policy. By incorporating this term into the optimization process, our method explicitly accounts for the causal estimated treatment effect difference between the target policy and the behavior policy. This causal perspective enhances the agent’s ability to reason about the differences between the target policy and the behavior policy used to generate the experiences, thereby mitigating the fundamental divergence between on-policy and off-policy learning. Importantly, our method reuses existing data of the behavior policy network values that are usually discarded to improve sample efficiency at a negligible cost.
We evaluate our proposed method, SUFT [^2] We name this novel theoretical result the SUFT causal bound, representing the initials of Sample efficient, Upper bound, Factual loss, and Treatment effect., on existing off-policy agents such as DQN and SAC, as well as on-policy agents like PPO, which benefit from our method due to the deviation between behavior and target policies during optimization. Throughout comprehensive experiments on the Atari 2600 and MuJoCo domains, agents using the SUFT OPE term outperform the same ones without it. Notably, we achieve up to 383% higher reward ratio, as shown in Figure˜2, indicating an improvement in the training convergence rate compared to an identical agent without the additional term. In addition, we reduce the experience replay buffer size by up to 96% while outperforming the baseline agent that uses a 25 times larger buffer by up to 437.05%. These significant improvements highlight the potential of our method to make DRL more computationally efficient and accessible.
Figure 2: Mean reward ratio comparison between agents using the additional SUFT OPE term and the baseline agents without it across 57 Atari games and five MuJoCo environments, highlighting the profound reward gains across diverse agents and domains.
2 Related Work
2.1 Causal Reinforcement Learning
Causal reinforcement learning (CRL) is an emerging subfield of reinforcement learning (RL). CRL seeks to enhance agents’ performance by incorporating causal inference methodologies into the learning process. The causal perspective allows the agent to learn cause-and-effect knowledge, resulting in more informed and effective decision-making [28, 29].
In causal inference, a treatment typically refers to performing a single action, commonly termed as performing an intervention [30, 31]. On the other hand, the majority of works in CRL define treatment as a sequence of actions determined by a dynamic treatment plan [32, 33, 34, 35]. In this work, we adopt the latter approach and consider the treatment to be the agent’s policy. This aligns with the RL framework, which focuses on evaluating the long-term effects of policies rather than the immediate impact of individual actions. Recent research in CRL covers a variety of areas, such as advancing generalizability [36], promoting safe agents [37, 38], enhancing sample efficiency [39, 40, 41], and addressing spurious correlations [42, 43, 44]. Our work falls in the field of CRL and primarily contributes to the latter two directions.
CRL for Addressing Spurious Correlations:
Spurious correlations in off-policy RL often arise as a result of learning based on experiences from an old behavior policy. These experiences might not accurately reflect the states and actions the target policy is likely to encounter, resulting in deceptive associations between actions and outcomes that limit the agent’s ability to make reliable decisions. Several works leverage OPE as a systematic way to evaluate the quality of a policy using off-policy data, which helps mitigate the impact of spurious correlations during policy optimization [45, 46].
CRL for Addressing Sample Inefficiency:
Causality provides a valuable perspective for creating sample-efficient agents, with research focusing on three key directions: representation learning [47, 48, 49], directed exploration [50, 51], and data augmentation [52, 53]. While many works improve sample efficiency by generating new data, our method utilizes existing discarded data from past experiences, achieving sample efficiency at a negligible cost. This novel strategy sets our work apart, demonstrating that DRL agents can achieve sample efficiency with marginal computational overhead.
2.2 Potential Outcomes Causal Bounds
The Neyman-Rubin potential outcomes framework [54] is a cornerstone of causal inference, providing a robust theoretical basis for estimating treatment effects. Several works have utilized this framework to develop causal bounds [55, 56]. For instance, Shalit et al. [57] establish a generalization-error bound to estimate individual treatment effects from observational data, addressing unobserved counterfactuals by formalizing a lemma that bounds this loss. Our work expands on this foundational framework and introduces several key changes.
First, while most works apply these bounds within general machine learning (ML) settings [58, 59], we redefine the terms to align with the DRL framework. Second, while other studies consider treatment as a single intervention [60, 61], we address treatments as a sequence of actions within a treatment plan. The primary and novel aspect of our work lies in reversing the traditional loss-bounding methodologies. Whereas existing methods typically bound the counterfactual loss [62, 63], our approach does the exact opposite and bounds the factual loss. This pivotal shift is feasible within the RL framework since off-policy methods use counterfactual data while factual data remains unobserved. Finally, whereas other approaches mainly focus on the scenario of a single target treatment and a single control treatment [64, 65], we scale this scenario to a single target treatment and $N$ control treatments. This multiple control treatments scenario allows us to leverage the replay buffer, which contains experiences generated from $N$ behavior policies, enabling effective evaluation of the target policy against a broader spectrum of previous behavior policies.
3 Preliminaries
In this section, we present some fundamental conceptual ideas necessary to understand our method.
Causal and Counterfactual Reasoning:
Causal reasoning refers to the process of identifying and understanding the relationships between causes and their effects. Counterfactual reasoning [66], a subfield of causal reasoning, goes a step further by imagining alternative scenarios to answer "what-if" questions, such as "What would have happened if a different treatment had been applied?" as Pearl & Mackenzie [67] elaborate in "The Book of Why". This form of reasoning enables the evaluation of potential outcomes conditioned under different treatments. In RL, incorporating counterfactual reasoning enhances agents’ capabilities to evaluate and optimize policies in complex, dynamic environments. By retrospectively analyzing alternative policies and their outcomes, agents can improve decision-making and policy optimization. In this paper, counterfactual reasoning is implemented through the integration of a causal bound into the agent’s loss function.
Off-Policy Evaluation:
Off-policy evaluation is a fundamental concept in RL [68], which aims to numerically evaluate the performance of a desired policy $\pi$ using historical data $D$ that was collected under different policies. As outlined in [69], the desired policy being optimized is referred to as the target policy, while the policies used to generate experiences are termed the behavior policies. OPE is particularly important in offline RL, where the agent learns the target policy solely from a static dataset $D$ without interacting with the environment. In contrast, this work focuses on applying OPE to online DRL agents, which continuously interact with the environment while updating the experience replay buffer $D$.
4 SUFT Causal Bound
In this section, we formalize our theoretical causal result and detail our method.
4.1 Rationale
To illustrate our method and provide a general intuition, we begin with a straightforward example of a greedy off-policy DQN agent. This agent observes the current state of the environment, uses its Q-network to compute Q-values, and selects the action with the highest Q-value for interacting with the environment. As the agent continuously interacts with the environment, it stores the experiences in a replay buffer and periodically updates its Q-network, which determines its target policy.
During Q-network updates, an ideal scenario would be to replace all experiences with fresh samples generated directly from the target policy without any additional cost. This would preserve the sample efficiency from the off-policy approach while resolving the divergence between the behavior and target policies. However, since such a scenario is impractical, we introduce a novel mathematical result and adapt it to the DRL framework to bridge between on-policy and off-policy methods.
Within the causal framework, the on-policy loss corresponds to the factual loss when the treatment being optimized matches the treatment applied for generating the observations. This is analogous to optimizing the target policy using experiences from an identical behavior policy. Similarly, the off-policy loss aligns with the counterfactual loss when the treatment being optimized differs from the one used to generate the observations. This is equivalent to optimizing the target policy using experiences from a different behavior policy.
In this work, we establish a causal upper bound on the unobserved factual loss using the observed counterfactual loss and an additional term, which is the estimated treatment effect of the target treatment compared to $N$ control treatments. After formalizing the upper bound within the causal framework, we seamlessly integrate it into the DRL framework as an upper bound for the on-policy loss using the standard off-policy loss and the estimated treatment effect, which serves as an OPE.
Throughout this paper, we demonstrate the causal definitions alongside corresponding DQN definitions to provide a clear and intuitive representation of the reduction from the causal framework to the DRL framework. Nevertheless, our method is applicable to any DRL agent with a V or Q-value network that influences its policy, such as DQN and Actor-Critic architectures, as long as a divergence exists between the target and behavior policies.
4.2 Setup and Definitions
Here, we present the necessary mathematical definitions required to formalize our causal upper bound. Comprehensive definitions and proofs, along with a detailed reduction from the causal framework to the DRL framework, are provided in Appendix˜B.
To simplify the equations, we present the binary treatment case, where there is a single target treatment, denoted by $t=1$, and a single control treatment, denoted by $t=0$. In practical DRL applications, we extend the binary case to a multiple control treatment case involving a single target policy and $N$ behavior policies, reflecting the existence of experiences generated from different behavior policies in the replay buffer. The proofs and definitions for the multiple control treatment case are in Appendix˜B.
We employ the Neyman-Rubin potential outcomes framework and define the following: $\mathcal{X}\subset\mathbb{R}^{d}$ is the observation space. $\mathcal{Y}\subset\mathbb{R}$ is the outcome space. $\mathcal{T}=\{0,1\}$ is the binary treatment space. Under this framework, there are two potential outcomes depending on the treatment assignment: if $t=1$, the observed outcome is $y=Y_{1}$, and if $t=0$, the observed outcome is $y=Y_{0}$. We assume there exists a joint distribution $P(X,T,Y_{0},Y_{1})$ with the following components: $q_{t}:=\Pr(T=t)$ is the treatment probability. $P^{t}_{X}:=P(X|T=t)$ is the conditional distribution of observations. $P^{x}_{Y_{1}}:=P(Y_{1}|X=x)$ is the conditional distribution of potential outcomes given the target treatment, and $P^{x}_{Y_{0}}:=P(Y_{0}|X=x)$ is the conditional distribution of potential outcomes given the control treatment. We also assume that there are no hidden confounders in the observational data, which makes the estimated treatment effect identifiable. This is formalized by assuming that the standard strong ignorability condition holds: $(Y_{0},Y_{1})\mathrel{\perp\!\!\!\perp}T|X$, and $0<p(t=1|x)<1$ for all $x$. Strong ignorability is a sufficient condition for the estimated treatment effect to be identifiable [70, 57]. To integrate the potential outcomes framework into the ML context, we define a neural network $\phi(x;\theta_{t}):\mathcal{X}\to\mathcal{Y}$, parameterized by weights $\theta_{t}$ corresponding to the performed treatment $t\in\mathcal{T}$.
We now present the analogous definitions within the DRL framework for a DQN agent, assuming the DRL framework is defined as a Markov decision process (MDP).
$\mathcal{X}:=(\mathcal{S\times A})$ is the state-action space. Potential outcome $y\sim P^{(s,a)}_{Y_{t}}$ is the accumulated reward over $n$ time steps, starting from state $s$, taking the initial action $a$, and following policy $\pi_{t}$. The hypothesis $\phi(x;\theta_{t}):=Q(s,a;\theta_{t}):(\mathcal{S\times A})\to\mathcal{Y}$ is the Q-values network, which influences the policy $\pi_{t}$.
Let $\mathbf{L}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$ represent a loss function. We rely on the following inequality assumption, which is a variation of the inverse triangle inequality:
Assumption 4.1.
$\mathbf{L}(x,y)-\mathbf{L}(x^{\prime},y^{\prime})$ $\leq\mathbf{L}(x,x^{\prime})+\mathbf{L}(y,y^{\prime}).$
The proof for the $L_{1}$} loss is in Appendix˜B [^3] DRL agents are often trained using $L_{2}$ loss, for which 4.1 does not hold. While the assumption is needed to prove Theorem 4.10, similar experimental results were obtained with both $L_{1}$ and $L_{2}$ losses. To provide intuition for the observed $L_{2}$} performance, we empirically examined the validity of 4.1 using $L_{2}$. This analysis is detailed in Appendix D..
Now, we define the expected factual and counterfactual losses, which are used to formulate the bound.
Definition 4.2.
The expected outcome loss:
$\ell_{\phi}(x,t)$ $:=\mathbb{E}_{y\sim P^{x}_{Y_{t}}}[\mathbf{L}(y,\phi(x;\theta_{t}))].$
The corresponding definition in DQN:
Definition 4.3.
The expected temporal difference loss:
$\ell_{Q}(s,a,t)$ $:=\mathbb{E}_{y\sim P^{(s,a)}_{Y_{t}}}[\mathbf{L}\left(y,Q(s,a;\theta_{t})\right)].$
Note that in the DRL framework, the expected outcome loss refers to the temporal difference loss between the accumulated rewards and the Q-value estimation.
Definition 4.4.
The expected factual outcome loss:
$\epsilon^{1}_{F_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,1)].$
$\epsilon^{0}_{F_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,0)].$
$\epsilon_{F_{\phi}}$ $:=q_{1}\epsilon^{1}_{F_{\phi}}+q_{0}\epsilon^{0}_{F_{\phi}}.$
To formalize the corresponding definition in the DRL framework, consider a replay buffer $D_{1}$ containing samples produced by a behavior policy that matches the target policy. In this context, the on-policy loss is a special case of the expected factual loss, and for DQN, it is defined as:
Definition 4.5.
The expected on-policy loss:
$\epsilon_{\text{On-Policy}_{Q}}$ $:=\mathbb{E}_{\begin{subarray}{c}(s,a)\sim D_{1}\end{subarray}}[\ell_{Q}(s,a,t_{\text{target}})].$
Note that this loss is the temporal difference loss with experiences collected on-policy, with identical target and behavior policies from $D_{1}$.
Definition 4.6.
The expected counterfactual outcome loss:
$\epsilon^{1}_{CF_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,1)].$
$\epsilon^{0}_{CF_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,0)].$
$\epsilon_{CF_{\phi}}$ $:=q_{0}\epsilon^{1}_{CF_{\phi}}+q_{1}\epsilon^{0}_{CF_{\phi}}.$
To establish the corresponding definition within the DRL framework, consider a replay buffer $D_{0}$ that stores experiences generated by a behavior policy that differs from the target policy. In this context, the off-policy loss is a special case of the expected counterfactual loss, and for DQN, it is defined as:
Definition 4.7.
The expected off-policy loss:
$\epsilon_{\text{Off-Policy}_{Q}}$ $:=\mathbb{E}_{\begin{subarray}{c}(s,a)\sim D_{0}\end{subarray}}[\ell_{Q}(s,a,t_{\text{target}})].$
Note that this loss is the temporal difference loss with experiences gathered off-policy, with different target and behavior policies from $D_{0}$.
Additionally, in DRL applications, we consider the scenario of a replay buffer containing experiences generated from $N$ distinct behavior policies. The definitions for this case are in Appendix˜B.
We now define our method’s additional estimated treatment effect term in both the causal and the DRL frameworks:
Definition 4.8.
The estimated treatment effect loss:
$\psi_{\phi}^{1}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{0}),\phi(x;\theta_{1}))].$
$\psi_{\phi}^{0}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{1}),\phi(x;\theta_{0}))].$
$\psi_{\phi}$ $:=q_{0}\psi_{\phi}^{1}+q_{1}\psi_{\phi}^{0}.$
The interpretation for the estimated treatment effect in DRL is an OPE, and for DQN, it is defined as:
Definition 4.9.
The SUFT OPE term:
$\psi_{\text{SUFT}_{Q}}$ $:=\mathbb{E}_{(s,a)\sim D_{0}}[\mathbf{L}\left(Q(s,a;\theta_{\text{behavior}}),Q(s,a;\theta_{\text{target}})\right)].$
This term aligns with the principles of OPE, focusing on quantifying the discrepancy between the target policy and the behavior policy using off-policy data. Once the data is stored in the buffer, it is used as observational data.
4.3 Causal Bound Theorem
Having established the key definitions, we now present our novel factual loss upper bound theorem:
Theorem 4.10.
The expected factual outcome loss, $\epsilon_{F_{\phi}}$, is bounded by the expected counterfactual outcome loss, $\epsilon_{CF_{\phi}}$, the estimated treatment effect loss, $\psi_{\phi}$, and a constant term $\delta$, independent of $\phi$:
$\epsilon_{F_{\phi}}$ $\leq\epsilon_{CF_{\phi}}+\psi_{\phi}+\delta.$
The proof appears in Appendix˜B.
The interpretation of Theorem˜4.10 in the DRL framework establishes an upper bound to the on-policy loss, by the off-policy loss, the SUFT OPE term, and a constant term $\delta$. A certain difference exists in the observational data between the causal and the DRL framework, which is detailed in Appendix˜B.
$\epsilon_{\text{On-Policy}_{Q}}$ $\leq\epsilon_{\text{Off-Policy}_{Q}}+\psi_{\text{SUFT}_{Q}}+\delta.$
Note that since $\delta$ is independent of $Q$, it can be omitted during the gradient-based optimization of $Q$.
To the best of our knowledge, we are the first to introduce such an upper bound within the causal framework on the factual loss, adapting it to the DRL framework, thereby bounding the on-policy loss. Importantly, based on our causal bound theorem, the adaptation to the DRL framework is not constrained to DQN agents but can be applied to every off-policy agent with V or Q-value networks that influence its policy. This scenario is thoroughly detailed in Appendix˜B. Furthermore, our method is also applicable to on-policy agents, such as PPO, which benefits from it due to the divergence between the behavior and target policies that arises during epoch optimization.
4.4 Recycling Value Network Outputs
To compute the SUFT OPE term in the causal bound, we need to calculate the Q-network values corresponding to both the target and the behavior policies. While the Q-values for the target policy can be easily computed using the current Q-network, the Q-network values for the old behavior policies are not addressable since storing the network weights for every behavior policy is impractical. Notably, during the agent’s training process, the Q-values matching the behavior policies are already calculated when they are used to select the action to perform. These values are calculated for the purpose of exploitation and then discarded once the action is executed. The novel contribution of our approach lies in utilizing and reusing the value network outputs in order to compute the SUFT OPE term. This allows the reuse of data that is overlooked while enhancing the loss optimization with significant causal insights in the form of the estimated treatment effect. Our approach effectively transforms overlooked data into valuable causal insights, metaphorically turning sand data into gold.
To obtain value network outputs in practice, we store them in the experience replay buffer. For agents employing a Q-network, only a single Q-value corresponding to the behavior policy is stored, resulting in an experience tuple of the form $\left(s,a,r,s^{\prime},Q(s,a;\theta_{\text{behavior}})\right)$, instead of the standard $\left(s,a,r,s^{\prime}\right)$. For agents using a V-network, the state value output is stored instead, yielding an experience tuple of the form $\left(s,a,r,s^{\prime},V(s;\theta_{\text{behavior}})\right)$, instead of the standard $\left(s,a,r,s^{\prime}\right)$.
4.5 Universal Implementation of the SUFT OPE Term
Following the formalization of the SUFT causal bound, in Section˜4.3, we adopt the well-established approach of optimizing a bound on the loss, a technique commonly employed in algorithms like expectation–maximization [71] and variational inference [72]. Consequently, rather than directly optimizing the on-policy loss, we focus on optimizing the SUFT causal bound. In addition, to improve the flexibility and accuracy of the loss optimization, we incorporate a $\lambda_{TF}$ coefficient into the SUFT OPE additional term. In the case of DQN, the SUFT causal bound objective term is as follows:
$\epsilon_{\text{Off-Policy}_{Q}}$ $+\lambda_{\text{TF}}\cdot\psi_{\text{SUFT}_{Q}}.$
In value-based methods such as DQN agents, the SUFT OPE term is seamlessly added to the standard loss function, while in Actor-Critic architectures, it is incorporated into the critic’s loss. We provide a detailed pseudocode of our method algorithms for both DQN and PPO in Appendix˜C.
5 Experiments
In this section, we provide a comprehensive empirical analysis of our proposed method, emphasizing the impact of adding the SUFT OPE term to the agent’s standard loss function across a broad and diverse range of environments. In this section, we present results obtained using $L_{2}$} loss, allowing us to compare our method to standard DRL agents. The results for $L_{1}$} loss are detailed in Appendix˜D.
5.1 Experimental Setup
We assess our method’s performance across two distinct domains: The first consists of 57 Atari games with a discrete action space, which contain a range of decision-making scenarios, including immediate versus long-term strategies, varying image background complexities, and dynamic on-screen objects. The second includes five MuJoCo environments that simulate robotic motions with a continuous action space that provides a robust platform for dynamic physical interactions.
In the Atari domain, we perform a wide range of experiments using DQN [21], which serves as the foundational algorithm for numerous Q-value network agents, Vanilla DQN, and Double DQN [73], as well as PPO [20], which is a widely adopted agent that has become a standard in the field. The Vanilla DQN results are detailed in Appendix˜D. For the MuJoCo domain, we evaluate our method using SAC [22], a continuous control off-policy actor-critic agent known for its robust performance, and with PPO. Across these settings, we explore different values of $\lambda_{TF}$, resulting in more than 10K individual experiments. This extensive and versatile experimentation ensures statistically significant findings, providing confidence in the robustness of the results and enabling us to draw more reliable conclusions about our method’s performance. More details about tuning $\lambda_{TF}$ are in Appendix˜D.
Training Details:
To evaluate SUFT’s sample efficiency, we conduct experiments with constrained environment interactions. For the off-policy agents, we use a relatively small experience replay buffer size of 4K. Each agent configuration is independently trained 10 times per environment, with the training process limited to 400K steps in the Atari domain and 1M steps in the MuJoCo domain. Performances are analyzed by taking the median smoothed training reward from the 10-trial set, ensuring a more reliable evaluation while mitigating the influence of randomness on the results. This constrained resources experimentation is designed to demonstrate our method’s sample efficiency, and it is a common technique in DRL. [74] Results for a larger buffer size are in Appendix˜D.
Implementation Details:
We use the Gymnasium environments API [75] to train on the Atari [76] (version v5) and MuJoCo [77] (version v4) domains. For training the agents, we are using the default architecture and hyperparameter configurations as implemented in the "Stable-Baselines3" DRL library [78]. In the Atari games, we apply standard preprocessing, which includes resizing observations to 84x84 grayscale images and stacking four consecutive frames. This preprocessing is implemented using the library’s default Atari wrapper.
Figure 3: Learning curves comparison between agents using the SUFT OPE term and the baseline agents without it across selected Atari games. The red line indicates human-level performance, showing that SUFT not only surpasses the baseline agent but can even exceed human rewards. Top: Double DQN; Bottom: PPO.
5.2 Controlled Comparison with Baseline Agents
We evaluate the impact of incorporating the SUFT OPE term by comparing agents that include our proposed term to baseline agents without it, ensuring both types share identical architecture and hyperparameters, where the presence of our proposed term is the only difference between them. This controlled experiment is designed to isolate the influence of our method, ensuring fair comparisons.
In addition, we conduct a $t$-test for each environment against a two-sided alternative using 10 different random seed runs, where the null hypothesis is the baseline result. The $p$-values provide robust statistical significance results for evaluating the effectiveness of our proposed method.
The complete experimental results, the evaluation formulas, and more empirical analysis of our method are detailed in Appendix˜D.
Atari:
On the Atari benchmark, SUFT demonstrates significant improvements for all of the agents. On the DQN, Double DQN, and PPO agents, SUFT surpasses the baseline agent in 52, 48, and 50 games out of the 57 games, respectively, with statistically significant gains ($p$-value << 0.05) in 49, 37, and 40 games.
We also analyze the results for the valid Atari games (according to the evaluation formula in Appendix˜D). For the DQN agent, SUFT outperforms the baseline agent in 28 out of the 29 valid games, achieving a reward improvement of over 100% in 48.3% of them and over 10% in 93.1% of them, as shown in Figure˜11. For the Double DQN agent, SUFT outperforms the baseline agent in 35 out of the 40 valid games, achieving a reward improvement of over 100% in 20% of them and over 10% in 72.5% of them, as shown in Figure˜1 (Left). For the PPO agent, SUFT surpasses the baseline agent in 39 out of the 42 valid games, achieving a reward improvement of over 100% in 38.1% of them and over 10% in 83.3% of them, as shown in Figure˜1 (Right).
Additionally, when evaluating the mean reward ratio, SUFT demonstrates profound gains relative to the baseline agents. For the DQN, Vanilla DQN, Double DQN, and PPO agents, SUFT achieves a mean reward ratio of 383.19%, 228.12%, 226.91%, and 197.34%, respectively, as shown in Figure˜2.
Moreover, in several Atari games, SUFT not only surpasses the baseline agent but also achieves higher rewards than an average human, as shown in Figure˜3. This achievement highlights the potential of our method to boost agents’ performance even above the human level, while keeping constrained environment interactions and resource demands. The human rewards are sourced from [79].
MuJoCo:
For the MuJoCo benchmark, SUFT demonstrates improvements in the SAC and PPO agents. For the SAC agent, SUFT surpasses the baseline agent in 4 out of 5 environments, with statistically significant gains in 3 environments. For the PPO agent, SUFT surpasses the baseline agent in 4 out of 5 environments (in the other environment, the score is the same), with statistically significant gains in 2 environments. With regard to the mean reward ratio, SUFT achieves 224.53% for the SAC agent and 111.21% for the PPO agent, as shown in Figure˜2.
5.3 Buffer Size Reduction
We also evaluate the sample efficiency of SUFT by comparing its performance to an identical DQN agent utilizing a larger 100K experience replay buffer size. Remarkably, SUFT achieves a mean reward ratio gain of 437.05%, outperforming an identical agent that uses a 25 times larger buffer size.
These results underscore the potential of our approach to obtain high-performance DRL agents with significantly smaller buffers, paving the way for more computationally efficient and accessible DRL methods by enhancing sample efficiency at a negligible cost.
5.4 Ablation Study
We conduct an ablation study to isolate the contribution of the SUFT OPE term to DQN performance. As shown Table˜1, we compare SUFT DQN against several alternatives: increasing the buffer size, a Vanilla DQN, and a Double DQN. While enlarging the buffer or adopting more complex DQN variants leads to improvements, incorporating the SUFT OPE term yields the highest Human Normalized Mean of 63.45%, outperforming all other baselines that require higher resources and computational costs. These results highlight the potential of our method to improve the agent’s performance by enhancing sample efficiency at a negligible cost, providing a simple yet highly effective method for improving agent performance at a negligible cost.
Table 1: An ablation study on the DQN agent comparing the SUFT additional term impact against enlarging the buffer size, adding a Vanilla DQN, and adding a Double DQN. The DQN, Vanilla DQN, Double DQN, and SUFT DQN use a 4K buffer size. All agents are using $L_{2}$} loss.
Agent Human Normalized Mean (%)
DQN 25.04%
DQN, buffer 100K 29.29%
Vanilla DQN 41.26%
Double DQN 42.31%
SUFT DQN 63.45%
5.5 Causal Bound Optimization vs Inertia Regularizer
To isolate the effect of our causal mechanism and ensure that the significant improvements of our method stem from the causal bound optimization rather than merely from a regularization effect, we conduct a controlled experiment. Specifically, we evaluate a double DQN agent that adds an inertia regularizer to the standard temporal difference loss function in the form of:
$\left(Q_{\text{current network}}-Q_{\text{target network}}\right)^{2}$
Where $Q_{\text{target network}}$ denotes the slowly updated target network, as done in a standard double DQN. The experiment uses the same configurations as both the SUFT double DQN and the baseline double DQN, with the additional regularization term being the only difference from the baseline agent. The results in Table˜11 show that the inertia regularization significantly degrades double DQN performance. Resulting in an 83.21% mean reward ratio. This provides evidence that the gains from our method, which result in a 226.91% mean reward ratio improvement, arise from the causal bound optimization, where the old Q-network values used for the OPE term correspond to the behavior policy that generated the experiences, thereby mitigating the divergence between the on-policy and off-policy methods. These results indicate that our method functions not merely as a stability-promoting regularizer, but as a principled causal bound optimization.
6 Conclusion
This paper presents SUFT, a causal upper-bound loss optimization method that enhances DRL sample efficiency and reduces computational demands at a negligible cost. We begin by establishing a provable bound on the factual loss within the Neyman-Rubin potential outcomes framework and seamlessly adapting it to the DRL framework, showing that on-policy loss is bounded by every agent’s standard off-policy loss and an additional SUFT OPE term. This term is computable by storing past value network outputs in the experience replay buffer, thereby reusing discarded data to improve agents’ performance. Experimental results across diverse environments and agents show that our method leads to significant improvements in convergence rates and reward outcomes. Looking ahead to future work, we see physical robotics as a particularly promising domain, where SUFT’s reduced-resource method could make DRL more practical under real-world constraints.
Appendix A Causal Bound Notation Summary
$\mathbf{L(\cdot,\cdot)}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}:$ loss function.
$q_{t}:$ the probability of treatment $t$.
$P^{t}_{X}:$ the conditional distribution of observations.
$P^{x}_{Y_{1}}:$ the conditional distribution of potential outcomes given the target treatment.
$P^{x}_{Y_{0}}:$ the conditional distribution of potential outcomes given the control treatment.
$\phi(x;\theta_{t}):\mathcal{X}\to\mathcal{Y}:$ neural network hypothesis parameterized by weights $\theta_{t}$.
$\ell_{\phi}(x,t):$ the expected outcome loss.
$\epsilon_{F_{\phi}}:$ the expected factual outcome loss.
$\epsilon_{CF_{\phi}}:$ the expected counterfactual outcome loss.
$\psi_{\phi}:$ The estimated treatment effect loss.
$\delta:$ a constant term independent of $\phi$.
Appendix B Theoretical Result
B.1 Definitions and Proofs
In this section, we present the complete definitions and proofs to establish our theoretical result within the causal framework. Let $\mathbf{L}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$ represent a loss function.
Assumption B.1.
$\mathbf{L}(x,y)-\mathbf{L}(x^{\prime},y^{\prime})$ $\leq\mathbf{L}(x,x^{\prime})+\mathbf{L}(y,y^{\prime}).$
Proposition B.2.
˜B.1 holds for $L_{1}$} loss.
Proof.
$\mathbf{L}(x,y)-\mathbf{L}(x^{\prime},y^{\prime})$ $\leq|\mathbf{L}(x,y)-\mathbf{L}(x^{\prime},y^{\prime})|$ (1)
$=||x-y|-|x^{\prime}-y^{\prime}||$ (2)
$\leq|x-y-(x^{\prime}-y^{\prime})|$ (3)
$=|x-x^{\prime}+y^{\prime}-y)|$
$\leq|x-x^{\prime}|+|y^{\prime}-y|$ (4)
$=\mathbf{L}(x,x^{\prime})+\mathbf{L}(y,y^{\prime})$ (5)
Inequality 1 is by $x\leq|x|$, equality 2 is by the definition of $L_{1}$} loss, inequality 3 is by reverse triangle inequality $||x_{1}|-|x_{2}||\leq|x_{1}-x_{2}|$, inequality 4 is by triangle inequality $|x_{1}+x_{2}|\leq|x_{1}|+|x_{2}|$, equality 5 is by absolute symmetry |a|=|a||a|=|-a|, and the definition of $L_{1}$} loss. ∎
We start with the binary treatment case and employ the Neyman-Rubin potential outcomes framework to define the following: $\mathcal{X}\subset\mathbb{R}^{d}$ is the observation space. $\mathcal{Y}\subset\mathbb{R}$ is the outcome space. $\mathcal{T}=\{0,1\}$ is the binary treatment space. Under this framework, there are two potential outcomes depending on the treatment assignment: if $t=1$, the observed outcome is $y=Y_{1}$, and if $t=0$, the observed outcome is $y=Y_{0}$.
We assume a joint distribution $P(X,T,Y_{0},Y_{1})$ with the following components: $q_{t}:=\Pr(T=t)$ is the treatment probability. $P^{t}_{X}:=P(X|T=t)$ is the conditional distribution of observations. $P^{x}_{Y_{1}}:=P(Y_{1}|X=x)$ is the conditional distribution of potential outcomes given the target treatment, and $P^{x}_{Y_{0}}:=P(Y_{0}|X=x)$ is the conditional distribution of potential outcomes given the control treatment. We also assume that there are no hidden confounders in the observational data, which makes the estimated treatment effect identifiable. This is formalized by assuming that the standard strong ignorability condition holds: $(Y_{0},Y_{1})\mathrel{\perp\!\!\!\perp}T|X$, and $0<p(t=1|x)<1$ for all $x$. Strong ignorability is a sufficient condition for the estimated treatment effect to be identifiable [70, 57].
We now present the definitions used to formulate the causal upper bound.
Let $t\in\mathcal{T}$ be a treatment.
Definition B.3.
$\phi(x;\theta_{t}):\mathcal{X}\to\mathcal{Y}$ is a neural network hypothesis parameterized by weights $\theta_{t}$ corresponding to the performed treatment $t$.
Definition B.4.
The expected outcome loss:
$\ell_{\phi}(x,t)$ $:=\mathbb{E}_{y\sim P^{x}_{Y_{t}}}[\mathbf{L}(y,\phi(x;\theta_{t}))].$
Definition B.5.
The expected factual outcome loss:
$\epsilon^{1}_{F_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,1)].$
$\epsilon^{0}_{F_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,0)].$
Definition B.6.
The expected combined factual outcome loss:
$\epsilon_{F_{\phi}}$ $:=q_{1}\epsilon^{1}_{F_{\phi}}+q_{0}\epsilon^{0}_{F_{\phi}}.$
Definition B.7.
The expected counterfactual outcome loss:
$\epsilon^{1}_{CF_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,1)].$
$\epsilon^{0}_{CF_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,0)].$
Definition B.8.
The expected combined counterfactual outcome loss:
$\epsilon_{CF_{\phi}}$ $:=q_{0}\epsilon^{1}_{CF_{\phi}}+q_{1}\epsilon^{0}_{CF_{\phi}}.$
Definition B.9.
The estimated treatment effect loss:
$\psi_{\phi}^{1}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{0}),\phi(x;\theta_{1}))].$
$\psi_{\phi}^{0}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{1}),\phi(x;\theta_{0}))].$
Definition B.10.
The combined estimated treatment effect loss:
$\psi_{\phi}$ $:=q_{0}\psi_{\phi}^{1}+q_{1}\psi_{\phi}^{0}.$
B.2 Causal Bound Theorem
Lemma B.11.
$\mathbb{E}_{y_{1}\sim P^{x}_{Y_{1}}}[\mathbb{E}_{y_{0}\sim P^{x}_{Y_{0}}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]]$ $=\mathbb{E}_{y_{1}\sim P^{x}_{Y_{1}}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))].$
Similarly, for the opposite case:
$\mathbb{E}_{y_{0}\sim P^{x}_{Y_{0}}}[\mathbb{E}_{y_{1}\sim P^{x}_{Y_{1}}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))]]$ $=\mathbb{E}_{y_{0}\sim P^{x}_{Y_{0}}}[\mathbf{L}(y_{0},\phi(x;\theta_{0})))].$
Proof.
$\mathbb{E}_{y_{1}\sim P^{x}_{Y_{1}}}[\mathbb{E}_{y_{0}\sim P^{x}_{Y_{0}}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]]$ $=\int_{\mathcal{Y}}\left(\int_{\mathcal{Y}}\mathbf{L}(y_{1},\phi(x;\theta_{1}))\cdot p(y_{0}|x)dy_{0}\right)\cdot p(y_{1}|x)dy_{1}$
$=\int_{\mathcal{Y}}\mathbf{L}(y_{1},\phi(x;\theta_{1}))\cdot\left(\int_{\mathcal{Y}}p(y_{0}|x)dy_{0}\right)\cdot p(y_{1}|x)dy_{1}$
$=\int_{\mathcal{Y}}\mathbf{L}(y_{1},\phi(x;\theta_{1}))\cdot p(y_{1}|x)dy_{1}$
$=\mathbb{E}_{y_{1}\sim P^{x}_{Y_{1}}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]$
The proof for the opposite case is the same, replacing $y_{0}$ and $y_{1}$. ∎
Theorem B.12.
The expected factual outcome loss, $\epsilon_{F_{\phi}}$, is bounded by the expected counterfactual outcome loss, $\epsilon_{CF_{\phi}}$, the estimated treatment effect loss, $\psi_{\phi}$, and a constant term $\delta$, independent of $\phi$:
$\epsilon_{F_{\phi}}$ $\leq\epsilon_{CF_{\phi}}+\psi_{\phi}+\delta.$
Where $\delta$ is the following:
$\delta^{1}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{0},y_{1})]]].$
$\delta^{0}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{1},y_{0})]]].$
$\delta$ $:=q_{0}\delta^{1}+q_{1}\delta^{0}.$
Proof.
We begin by placing the loss definitions:
$\epsilon_{F_{\phi}}-\epsilon_{CF_{\phi}}$ $=q_{1}\epsilon^{1}_{F_{\phi}}+q_{0}\epsilon^{0}_{F_{\phi}}-q_{0}\epsilon^{1}_{CF_{\phi}}-q_{1}\epsilon^{0}_{CF_{\phi}}$ (6)
$=q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,1)]+q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,0)]$ (7)
$-q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\ell_{\phi}(x,1)]-q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,0)]$
$=q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]]+q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))]]$ (8)
$-q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]]-q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))]]$
$=q_{1}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]]-\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))]])$
$+q_{0}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))]]-\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]])$
Now we will use the linearity of expectation and Lemma˜B.11:
$\epsilon_{F_{\phi}}-\epsilon_{CF_{\phi}}$ $=q_{1}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]-\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))]])$ (9)
$+q_{0}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))]-\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))]])$
$=q_{1}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{1},\phi(x;\theta_{1}))-\mathbf{L}(y_{0},\phi(x;\theta_{0}))]]])$ (10)
$+q_{0}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))-\mathbf{L}(y_{1},\phi(x;\theta_{1}))]]])$
According to ˜B.1 we will get:
$\epsilon_{F_{\phi}}-\epsilon_{CF_{\phi}}$ $\leq q_{1}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{1},y_{0})+\mathbf{L}(\phi(x;\theta_{1}),\phi(x;\theta_{0}))]]])$ (11)
$+q_{0}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{0},y_{1})+\mathbf{L}(\phi(x;\theta_{0}),\phi(x;\theta_{1}))]]])$
$=q_{1}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{1},y_{0})]]]+\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{1}),\phi(x;\theta_{0}))])$ (12)
$+q_{0}(\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{0},y_{1})]]]+\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{0}),\phi(x;\theta_{1}))])$
$=q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{0}),\phi(x;\theta_{1}))]+q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{1}),\phi(x;\theta_{0}))]$
$+q_{0}\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{0},y_{1})]]]+q_{1}\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbf{L}(y_{1},y_{0})]]]$
$=q_{0}\psi_{\phi}^{1}+q_{1}\psi_{\phi}^{0}+q_{0}\delta^{1}+q_{1}\delta^{0}$ (13)
$=\psi_{\phi}+\delta$ (14)
Finally, we are getting the causal bound:
$\epsilon_{F_{\phi}}$ $\leq\epsilon_{CF_{\phi}}+\psi_{\phi}+\delta$
Equality 6 is by Definitions˜B.6 and B.8, equality 7 is by Definitions˜B.5 and B.7, equality 8 is by Definition˜B.4, equality 9 is by linearity of expectations, equality 10 is by linearity of expectations and by Lemma˜B.11, inequality 11 is by ˜B.1, equality 12 is by linearity of expectations and because $\phi$ is a constant on the expectation of $Y_{0},Y_{1}$, equality 13 is by Definition˜B.9 and $\delta^{1},\delta^{0}$ notation, equality 14 is by Definition˜B.10 and $\delta$ notation. ∎
B.3 Causal Bound for $N$ Control Treatments
We now adjust the definitions and Theorem˜B.12 proof from the binary treatment scenario to a single target treatment and N control treatments scenario.
We first set the definitions of the Neyman-Rubin potential outcome framework for the $N$ control treatments scenario: $\mathcal{T}=\{1,\ldots,N+1\}$ is the treatments space, where $t=j:\ 2\leq j\leq N+1$ is the control treatments, and $t=1$ is the target treatment compared against each control treatment. Under this framework, there are two types of potential outcomes depending on the treatment assignment: if $t=1$, the observed outcome $y=Y_{1}$ is from the target treatment, and $\forall j\in\{2,\ldots,N+1\}$ if $t=j$, the observed outcome $y=Y_{j}$ is from a control treatment.
We assume joint distribution $P(X,T,Y_{1},Y_{2},\ldots,Y_{N+1})$ with the following components: $q_{t}:=\Pr(T=t)$ is the treatment probability. $P^{t}_{X}:=P(X|T=t)$ is the conditional distribution of observations. $P^{x}_{Y_{1}}:=P(Y_{1}|X=x)$ is the conditional distribution of potential outcomes given the target treatment, and $\forall j\in\{2,\ldots,N+1\}$ $P^{x}_{Y_{j}}:=P(Y_{j}|X=x)$ is the conditional distribution of potential outcomes given a control treatment $t=j$.
Definition B.13.
The expected factual outcome loss for $N$ control treatments:
$\epsilon^{1}_{F_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,1)].$
$\forall j\in\{2,\ldots,N+1\}:$
$\epsilon^{j}_{F_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{j}_{X}\end{subarray}}[\ell_{\phi}(x,j)].$
Definition B.14.
The expected combined factual outcome loss for $N$ control treatments:
$\epsilon_{F_{\phi}}$ $:=q_{1}\epsilon^{1}_{F_{\phi}}+\sum_{j=2}^{N+1}q_{j}\epsilon^{j}_{F_{\phi}}.$
Definition B.15.
The expected counterfactual outcome loss for $N$ control treatments:
$\forall j\in\{2,\ldots,N+1\}:$
$\epsilon^{1,j}_{CF_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{j}_{X}\end{subarray}}[\ell_{\phi}(x,1)].$
$\epsilon^{j}_{CF_{\phi}}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\ell_{\phi}(x,j)].$
Definition B.16.
The expected combined counterfactual outcome loss for $N$ control treatments:
$\epsilon_{CF_{\phi}}$ $:=\sum_{j=2}^{N+1}\left(q_{j}\epsilon^{1,j}_{CF_{\phi}}+\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}\right).$
Definition B.17.
The estimated treatment effect loss for $N$ control treatments:
$\forall j\in\{2,\ldots,N+1\}:$
$\psi^{1,j}_{\phi}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{j}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{j}),\phi(x;\theta_{1}))].$
$\psi^{j}_{\phi}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbf{L}(\phi(x;\theta_{1}),\phi(x;\theta_{j}))].$
Definition B.18.
The combined estimated treatment effect loss for $N$ control treatments:
$\epsilon_{\psi_{\phi}}$ $:=\sum_{j=2}^{N+1}\left(q_{j}\psi^{1,j}_{\phi}+\frac{1}{N}q_{1}\psi^{j}_{\phi}\right).$
All the other definitions and proofs are the same as the binary case.
We now generalize Theorem˜B.12 from a single control treatment scenario to the $N$ control treatments scenario as follows:
Theorem B.19.
$q_{1}\epsilon^{1}_{F_{\phi}}+\sum_{j=2}^{N+1}q_{j}\epsilon^{j}_{F_{\phi}}$ $\leq\sum_{j=2}^{N+1}\left(q_{j}\epsilon^{1,j}_{CF_{\phi}}+\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}+q_{j}\psi^{1,j}_{\phi}+\frac{1}{N}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N}q_{1}\delta^{j}\right).$
Where $\delta$ is the following:
$\delta^{1,j}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{j}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{j}\sim P^{x}_{Y_{j}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{j},y_{1})]]].$
$\delta^{j}$ $:=\mathbb{E}_{\begin{subarray}{c}x\sim P^{1}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{j}\sim P^{x}_{Y_{j}}\end{subarray}}[\mathbf{L}(y_{1},y_{j})]]].$
$\delta$ $:=\sum_{j=2}^{N+1}\left(q_{j}\delta^{1,j}+\frac{1}{N}q_{1}\delta^{j}\right).$
Proof.
by induction.
Step 1: Base Case
For a single target treatment and a single control treatment, $N=1$, the following inequality holds:
$q_{1}\epsilon^{1}_{F_{\phi}}+q_{2}\epsilon^{2}_{F_{\phi}}-q_{2}\epsilon^{1,2}_{CF_{\phi}}-q_{1}\epsilon^{2}_{CF_{\phi}}$ $\leq q_{2}\psi^{1,2}_{\phi}+q_{1}\psi^{2}_{\phi}+q_{2}\delta^{1,2}+q_{1}\delta^{2}$
As we prove in Theorem˜B.12 where the control treatment is $t=2$ instead of $t=0$.
Step 2: Inductive Hypothesis
Assume that for $N-1\geq 2$ control treatments, the following inequality holds:
$\sum_{j=2}^{N}\left(\frac{1}{N-1}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N-1}q_{1}\epsilon^{j}_{CF_{\phi}}\right)$ $\leq$
$\sum_{j=2}^{N}\left(q_{j}\psi^{1,j}_{\phi}+\frac{1}{N-1}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N-1}q_{1}\delta^{j}\right)$
Step 3: Inductive Step
We will now prove the inequality for $N$ control treatments:
$\sum_{j=2}^{N+1}\left(\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}\right)$ $\leq\sum_{j=2}^{N+1}\left(q_{j}\psi^{1,j}_{\phi}+\frac{1}{N}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N}q_{1}\delta^{j}\right)$
Inductive Step proof.
$\sum_{j=2}^{N+1}\left(\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}\right)$ $=$
$\sum_{j=2}^{N}\left(\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}-q_{j}\epsilon^{1,j}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}\right)$ $+\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{N+1}\epsilon^{N+1}_{F_{\phi}}$
$-q_{N+1}\epsilon^{1,N+1}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{N+1}_{CF_{\phi}}$
$\leq\sum_{j=2}^{N}\left(q_{j}\psi^{1,j}_{\phi}+\frac{1}{N}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N}q_{1}\delta^{j}\right)$ (15)
$+\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{N+1}\epsilon^{N+1}_{F_{\phi}}$
$-q_{N+1}\epsilon^{1,N+1}_{CF_{\phi}}-\frac{1}{N}q_{1}\epsilon^{N+1}_{CF_{\phi}}$
$\leq\sum_{j=2}^{N}\left(q_{j}\psi^{1,j}_{\phi}+\frac{1}{N}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N}q_{1}\delta^{j}\right)$ (16)
$+q_{N+1}\psi^{1,N+1}_{\phi}+\frac{1}{N}q_{1}\psi^{N+1}_{\phi}$
$+q_{N+1}\delta^{1,N+1}+\frac{1}{N}q_{1}\delta^{N+1}$
$=\sum_{j=2}^{N+1}\left(q_{j}\psi^{1,j}_{\phi}+\frac{1}{N}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N}q_{1}\delta^{j}\right)$
We will get the following inequality:
$\sum_{j=2}^{N+1}\left(\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}+q_{j}\epsilon^{j}_{F_{\phi}}\right)$ $\leq\sum_{j=2}^{N+1}\left(q_{j}\epsilon^{1,j}_{CF_{\phi}}+\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}+q_{j}\psi^{1,j}_{\phi}+\frac{1}{N}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N}q_{1}\delta^{j}\right)$
Finally, we sum $\frac{1}{N}q_{1}\epsilon^{1}_{F_{\phi}}$ and get the causal bound for $N$ control treatments:
$q_{1}\epsilon^{1}_{F_{\phi}}+\sum_{j=2}^{N+1}q_{j}\epsilon^{j}_{F_{\phi}}$ $\leq\sum_{j=2}^{N+1}\left(q_{j}\epsilon^{1,j}_{CF_{\phi}}+\frac{1}{N}q_{1}\epsilon^{j}_{CF_{\phi}}+q_{j}\psi^{1,j}_{\phi}+\frac{1}{N}q_{1}\psi^{j}_{\phi}+q_{j}\delta^{1,j}+\frac{1}{N}q_{1}\delta^{j}\right)$
Inequality 15 is by the inductive hypothesis, Inequality 16 is from the proof of Theorem˜B.12 where the control treatment is $t=N+1$ instead of $t=0$. ∎
B.4 Reduction from Causal to DRL Framework
We now formalize a reduction from a causal framework to a DRL framework.
In the DRL framework, we consider the treatment to be the agent’s policy, where the target treatment matches the target policy, and the control treatments match the behavior policies.
Within the causal framework, the on-policy loss corresponds to the factual loss when the treatment being optimized matches the treatment applied for generating the observations. This is analogous to optimizing the target policy using experiences from an identical behavior policy. Similarly, the off-policy loss aligns with the counterfactual loss when the treatment being optimized differs from the one used to generate the observations. This is equivalent to optimizing the target policy using experiences from different behavior policies.
To formalize the corresponding definition in the DRL framework, consider a replay buffer $D_{1}$ containing samples produced by a behavior policy that matches the target policy. In this context, the on-policy loss is a special case of the expected factual loss.
Similarly, consider a replay buffer $D_{0}$ that stores experiences generated by behavior policies that differ from the target policy. In this context, the off-policy loss is a special case of the expected counterfactual loss.
Figure 4: Illustration diagram that demonstrates the reduction from the causal inference framework (Left) to the DRL framework (Right).
Q-Network Reduction:
The definitions for the scenario of an agent using a Q-value network:
$\mathcal{X}:=(\mathcal{S\times A})$ is the state-action space. Potential outcome $y\sim P^{(s,a)}_{Y_{t}}$ is the accumulated reward over n time steps, starting from state $s$, taking the initial action $a$, and following policy $\pi_{t}$. And the hypothesis $\phi(x;\theta_{t}):=Q(s,a;\theta_{t}):(\mathcal{S\times A})\to\mathcal{Y}$ is the Q-values network, which influences the policy $\pi_{t}$.
Definition˜B.4 in the DRL framework refers to the temporal difference loss between the accumulated rewards and the Q-value estimation, and it is defined as follows:
Definition B.20.
The expected temporal difference loss:
$\ell_{Q}(s,a,t)$ $:=\mathbb{E}_{y\sim P^{(s,a)}_{Y_{t}}}[\mathbf{L}\left(y,Q(s,a;\theta_{t})\right)].$
Definitions˜B.5 and B.13 in the DRL framework assigns the temporal difference loss with experiences collected on-policy, with identical target and behavior policies from $D_{1}$, and it is defined as follows:
Definition B.21.
The expected on-policy loss:
$\epsilon_{\text{On-Policy}_{Q}}$ $:=\mathbb{E}_{\begin{subarray}{c}(s,a)\sim D_{1}\end{subarray}}[\ell_{Q}(s,a,t_{\text{target}})].$
Definitions˜B.7 and B.15 in the DRL framework assigns the temporal difference loss with experiences gathered off-policy, with different target and behavior policies from $D_{0}$, and it is defined as follows:
Definition B.22.
The expected off-policy loss:
$\epsilon_{\text{Off-Policy}_{Q}}$ $:=\mathbb{E}_{\begin{subarray}{c}(s,a)\sim D_{0}\end{subarray}}[\ell_{Q}(s,a,t_{\text{target}})].$
Definitions˜B.9 and B.17 interpretation in the DRL framework is an OPE term, and it is defined as follows:
Definition B.23.
The SUFT OPE term:
$\psi_{\text{SUFT}_{Q}}$ $:=\mathbb{E}_{(s,a)\sim D_{0}}[\mathbf{L}\left(Q(s,a;\theta_{\text{behavior}}),Q(s,a;\theta_{\text{target}})\right)].$
Note that this term evaluates the target policy by measuring the loss between the Q-value estimations of the target policy and the behavior policies, given samples from $D_{0}$.
Theorems˜B.12 and B.19 interpretation in the DRL framework establishes an upper bound to the on-policy loss, combining the off-policy loss and the SUFT OPE term.
$\epsilon_{\text{On-Policy}_{Q}}$ $\leq\epsilon_{\text{Off-Policy}_{Q}}+\psi_{\text{SUFT}_{Q}}+\delta.$
Note that since the constant term, $\delta$, is independent of $Q$, it is disregarded during the gradient-based optimization of $Q$, so the SUFT causal bound objective term is as follows:
$\epsilon_{\text{Off-Policy}_{Q}}$ $+\lambda_{\text{TF}}\cdot\psi_{\text{SUFT}_{Q}}.$
V-Network Reduction:
Now we present the corresponding definitions for the scenario of an agent using a V network:
$\mathcal{X}:=\mathcal{S}$ is the state space. Potential outcome $y\sim P^{s}_{Y_{t}}$ is the accumulated reward over n time steps, starting from state $s$, and following policy $\pi_{t}$. And the hypothesis $\phi(x;\theta_{t}):=V(s;\theta_{t}):\mathcal{S}\to\mathcal{Y}$ is the V network, which influences the policy $\pi_{t}$.
Definition˜B.4 in the DRL framework refers to the temporal difference loss between the accumulated rewards and the V estimation, and it is defined as follows:
Definition B.24.
The expected temporal difference loss:
$\ell_{V}(s,t)$ $:=\mathbb{E}_{y\sim P^{s}_{Y_{t}}}[\mathbf{L}\left(y,V(s;\theta_{t})\right)].$
Definitions˜B.5 and B.13 in the DRL framework assigns the temporal difference loss with experiences collected on-policy, with identical target and behavior policies from $D_{1}$, and it is defined as follows:
Definition B.25.
The expected on-policy loss:
$\epsilon_{\text{On-Policy}_{V}}$ $:=\mathbb{E}_{\begin{subarray}{c}s\sim D_{1}\end{subarray}}[\ell_{V}(s,t_{\text{target}})].$
Definitions˜B.7 and B.15 in the DRL framework assigns the temporal difference loss with experiences gathered off-policy, with different target and behavior policies from $D_{0}$, and it is defined as follows:
Definition B.26.
The expected off-policy loss:
$\epsilon_{\text{Off-Policy}_{V}}$ $:=\mathbb{E}_{\begin{subarray}{c}s\sim D_{0}\end{subarray}}[\ell_{V}(s,t_{\text{target}})].$
Definitions˜B.9 and B.17 interpretation in the DRL framework is an OPE term, and it is defined as follows:
Definition B.27.
The SUFT OPE term:
$\psi_{\text{SUFT}_{V}}$ $:=\mathbb{E}_{s\sim D_{0}}[\mathbf{L}\left(V(s;\theta_{\text{behavior}}),V(s;\theta_{\text{target}})\right)].$
Note that this term evaluates the target policy by measuring the loss between the V estimations of the target policy and the behavior policies, given samples from $D_{0}$.
Theorems˜B.12 and B.19 interpretation in the DRL framework establishes an upper bound to the on-policy loss, combining the off-policy loss and the SUFT OPE term.
$\epsilon_{\text{On-Policy}_{V}}$ $\leq\epsilon_{\text{Off-Policy}_{V}}+\psi_{\text{SUFT}_{V}}+\delta.$
Note that since the constant term, $\delta$, is independent of $V$, it is disregarded during the gradient-based optimization of $V$. Thus, the SUFT causal bound objective term is as follows:
$\epsilon_{\text{Off-Policy}_{V}}$ $+\lambda_{\text{TF}}\cdot\psi_{\text{SUFT}_{V}}.$
Causal and DRL Framework Alignment:
An important alignment needs to be made to address a difference between the static observational data assumed by the causal framework and the dynamic experience replay buffer in the online DRL framework. In the causal framework, all observations are gathered once, so the data is fixed and irrevocably labeled either to be from the target treatment or from the control treatment in the binary treatment case. On the other hand, in the online DRL framework, the observational data changes over time, as we continuously populate the experience replay buffer and update the target policy. On each policy optimization step, all of the samples stored in the buffer that were labeled as the target policy are turned into samples that are labeled as a control policy. This phenomenon of changing the observational data labels when updating the policy in the DRL framework is a gap between the causal and DRL frameworks. In the on-policy case, the experience replay buffer contains only on-policy observations, meaning that $q_{1}=1$. When updating the target policy, this data becomes off-policy, and the data that was previously labeled as target policy turns into control policy, shifting the samples’ labels together with the treatment probability $q_{0_{\text{new}}}=q_{1_{\text{old}}}=1$.
The reduction we have made between the causal and the DRL framework above aligns the on-policy loss and off-policy loss terms with the corresponding terms in the causal framework. The on-policy loss is a special case of the factual loss where the observation data is filled with target treatment samples, corresponding to $\epsilon^{1}_{F}$ where $q_{1}=1$. The off-policy loss is a special case of the counterfactual loss where the observation data is filled with control treatment samples, corresponding to $\epsilon^{1}_{CF}$ where the labels and the treatment probability are shifted and $q_{0_{\text{new}}}=q_{1_{\text{old}}}=1$.
Appendix C SUFT Pseudocode
C.1 SUFT DQN Pseudocode
Algorithm 1 SUFT Deep Q-learning
Initialize replay memory $D_{0}$ to capacity NN
Initialize action-value function $Q$ with random weights $\theta$
for episode =1,M=1,M do
  Initialize start state $s_{1}$
  for $t=1,T$ do
   With probability $\rho$ select a random action $a_{t}$
   otherwise select $a_{t}=\arg\max_{a}Q\left(s_{t},a;\theta_{\text{behavior}}\right)$
   Execute action $a_{t}$ in the environment and observe reward $r_{t}$ and state $s_{t+1}$
   Store transition $\left(s_{t},a_{t},r_{t},s_{t+1},Q(s_{t},a_{t};\theta_{\text{behavior}})\right)$ in $D_{0}$
   Sample random mini-batch of transitions $\left(s_{j},a_{j},r_{j},s_{j+1},Q(s_{j},a_{j};\theta_{\text{behavior}})\right)$ from $D_{0}$
   Set $y_{j}=\begin{cases}r_{j}&\text{for terminal }s_{j+1}\\ r_{j}+\gamma\max_{a^{\prime}}Q(s_{j+1},a^{\prime};\theta_{\text{target}})&\text{for non-terminal }s_{j+1}\end{cases}$
   Perform a gradient descent step with respect to the network parameters $\theta_{\text{target}}$ on:
$\hat{\mathbb{E}}_{j}\left[\left(y_{j}-Q(s_{j},a_{j};\theta_{\text{target}})\right)^{2}+\lambda_{\text{TF}}\cdot\left(Q(s_{j},a_{j};\theta_{\text{behavior}})-Q(s_{j},a_{j};\theta_{\text{target}})\right)^{2}\right]$
  end for
end for
Our method extends the standard DQN algorithm [80] in two main ways: The first is by storing the old Q values $Q(s,a;\theta_{\text{behavior}})$ in the replay buffer. The second is by incorporating the additional SUFT OPE term into the objective function $\lambda_{\text{TF}}\cdot\left(Q(s_{j},a_{j};\theta_{\text{behavior}})-Q(s_{j},a_{j};\theta_{\text{target}})\right)^{2}$, which represents the estimated treatment effect. $\theta_{\text{behavior}}$ refers to the network weights used for action selection, the behavior policy, and $\theta_{\text{target}}$ represents the current network weights, which are optimized in the gradient descent step, the target policy.
C.2 SUFT PPO Pseudocode
Algorithm 2 SUFT Proximal Policy Optimization (PPO)
Initialize a rollout buffer $D_{1}$ with horizon $T$
Initialize policy function (actor) and value function (critic) with random weights $\theta$
for iteration $1,2,\ldots$ do
  for $t=1,T$ do
   Run policy $\pi_{\theta_{\text{behavior}}}(s_{t})$
   Store transition $\left(s_{t},a_{t},r_{t},\pi_{\theta_{\text{behavior}}}(a_{t}\mid s_{t}),V(s_{t};\theta_{\text{behavior}})\right)$ in $D_{1}$
  end for
  Compute advantage estimates $\hat{A}_{1},\ldots,\hat{A}_{t}$, and rewards to go $\hat{R}_{1},\ldots,\hat{R}_{t}$
  based on the stored values $V(s_{1};\theta_{\text{behavior}}),\ldots,V(s_{t};\theta_{\text{behavior}})$
  for $k=1,K$ epochs do
   Sample random mini-batch of transitions $\left(s_{j},a_{j},r_{j},\pi_{\theta_{\text{behavior}}}(a_{j}\mid s_{j}),V(s_{j};\theta_{\text{behavior}})\right)$
   from the rollout buffer $D_{1}$
   Compute probability ratio:
$r_{j}(\theta)=\frac{\pi_{\theta_{\text{target}}}(a_{j}\mid s_{j})}{\pi_{\theta_{\text{behavior}}}(a_{j}\mid s_{j})}$
   Optimize actor loss:
$\hat{\mathbb{E}}_{j}\left[\min\left(r_{j}(\theta)\hat{A}_{j},\;\text{clip}\left(r_{j}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{j}\right)\right]$
   Optimize critic loss:
$\hat{\mathbb{E}}_{j}\left[(V(s_{j};\theta_{\text{target}})-\hat{R}_{j})^{2}+\lambda_{\text{TF}}\cdot\left(V(s_{j};\theta_{\text{behavior}})-V(s_{j};\theta_{\text{target}})\right)^{2}\right]$
  end for
end for
Our method extends the standard PPO algorithm [20] by incorporating the additional SUFT OPE term into the PPO critic loss. Notably, PPO performs multiple epochs of optimization over the same rollout buffer $D_{1}$, leading to a divergence between the behavior and target policies. This divergence is the reason PPO benefits from our method, despite being an on-policy algorithm. $\theta_{\text{behavior}}$ refers to the network weights used to interact with the environment also referred to in [20] as $\theta_{\text{old}}$, the behavior policy, and $\theta_{\text{target}}$ represents the current network weights, which are optimized, also referred to in [20] as $\theta$, the target policy. During the first epoch, before optimizing the network, the behavior and target policies are identical, so the SUFT OPE term equals zero. After the first optimization step, the target policy diverges from the behavior policy, and the SUFT OPE term now quantifies the estimated treatment effect between the target and behavior policies. In addition, the $V(s_{t};\theta_{\text{behavior}})$ values are already calculated during the calculation of the advantage estimates $\hat{A}_{t}$, and rewards to go $\hat{R}_{t}$. Thus, our method effectively reuses these stored network outputs for computing the SUFT OPE term, similar to the approach used in our SUFT DQN implementation.
Appendix D Full Experimental Results
In this section, we present detailed and comprehensive experimental results for both the Atari and the MuJoCo domains, as well as the formulas and criteria used to evaluate the results.
D.1 Negligible Computational Overhead
We quantify the computational and space overhead introduced by computing the SUFT OPE term and storing old Q-values in the experience replay buffer. We run our experiments on a single NVIDIA L4 GPU, and for the Vanilla DQN agent, the additional computation takes approximately 23.77 seconds, compared to a total training time of 2,980 seconds. This corresponds to a 0.8% computational overhead. In terms of space, we append a single 64-bit scalar to each stored experience. For a 4K experience replay buffer, this results in an additional 0.0305 MB, compared to 972.85 MB without it. This is equivalent to a 0.003% space overhead.
D.2 Tuning the $\lambda_{TF}$ Coefficient
In order to evaluate the impact of the $\lambda_{TF}$ coefficient on improving the loss accuracy, we conducted experiments using various $\lambda_{TF}$ values. Adding the $\lambda_{TF}$ demonstrates a consistent improvement across a range of $\lambda_{TF}$ values. This indicates that precise hyperparameter tuning is unnecessary, and as long as the SUFT OPE term is added with a coefficient within a valid range of $\lambda_{TF}$, the loss optimization improves. The $\lambda_{TF}$ value can be easily tuned through random search hyperparameter optimization within a valid range. Specifically, we encounter that a sufficient range to search from is between $\lambda_{TF}=0.5$ and $\lambda_{TF}=1.5$. The $\lambda_{TF}$ values that yield the best performance are $\lambda_{TF}=1$ for all the DQN agents, $\lambda_{TF}=2.5$ for the PPO agent using the Using $L_{1}$} Loss and $\lambda_{TF}=5$ for the PPO agent using the Using $L_{2}$} Loss in the Atari domain. In the MuJoCo domain, SAC performs best at $\lambda_{TF}=0.6$ and PPO at $\lambda_{TF}=1.8$.
D.3 Evaluation Formulas
To evaluate our method, we train every agent-environment configuration for 10 different runs and report the median result for each setup. Since this is an even-numbered list, there are two central values, and we use the higher of the two as the median. The valid environments used for calculating the improvement percentage and mean reward ratio percentage formulas are those where the numerator and denominator are greater than one. This restriction allows us to exclude cases where the reward is only marginally higher than the random reward, which could otherwise yield arbitrarily large ratios. By doing so, we ensure that the metric remains stable and effectively reflects meaningful improvements.
Improvement Percentage Formula:
The improvement percentage for comparing our approach to an identical agent without the additional term is calculated with the following formula:
[…]
For better visualization of the results, we display them in log scale with the following formula:
$\log_{10}\left[\text{Improvement Percentage Formula}+1\right]$
We add 1 to the improvement percentage formula to avoid the log of 0 and fractions, which result in a negative log. This addition has no impact on the results.
Mean Reward Ratio Percentage Formula:
The mean reward ratio percentage for comparing our approach to an identical agent without the additional term is calculated with the following formula:
$\frac{1}{\mid\text{Valid Envs}\mid}\sum_{\forall\text{env}}\frac{\text{SUFT reward}\ \ \ \ \ -\text{Random reward}}{\text{Baseline reward }-\text{Random reward}}\cdot 100\%.$
Note that for the Atari domain, the human normalized results are calculated using the above formula, where the baseline reward is replaced with the human reward.
D.4 ˜4.1 Empirical Analysis with $L_{2}$
DRL agents are often trained using $L_{2}$ loss, for which ˜4.1 does not hold. To provide intuition for the observed $L_{2}$} performance, we empirically examine the validity of the $L_{2}$ loss for ˜4.1 using synthetic and domain-specific data. For the synthetic data, we generate one billion independent samples from the standard normal distribution, $x,y,x^{\prime},y^{\prime}\sim\mathcal{N}(0,1)$, and evaluate whether:
$\left(x-y\right)^{2}-\left(x^{\prime}-y^{\prime}\right)^{2}$ $\leq\left(x-x^{\prime}\right)^{2}+\left(y-y^{\prime}\right)^{2}$ (17)
This experiment corresponds directly to the inequality form of ˜4.1 when applied to the $L_{2}$ loss. We find that inequality 17 holds in approximately 82% of random samples, demonstrating that while ˜4.1 is not guaranteed theoretically, it is satisfied in the vast majority of cases. To further validate this observation in a domain-specific dataset, we conduct an experiment in the Atari 57 benchmark using a DQN agent trained for 400K steps with a batch size of 32. Since our method operates in an off-policy setting, the left side of the causal bound in Theorem˜4.10 is not computable from the observed data, as we don’t observe the on-policy experiences (this is the reason we optimize on the causal bound in our method). For this reason, we evaluate only inequality 11 used in the proof of Theorem˜B.12, considering the off-policy terms on both sides (with $q_{0}=1$):
$(\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))-\mathbf{L}(y_{1},\phi(x;\theta_{1}))]]])$ $\leq$ (18)
$(\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{0},y_{1})+\mathbf{L}(\phi(x;\theta_{0}),\phi(x;\theta_{1}))]]])$
Because the true values of $y_{0}$ and $y_{1}$ are not directly observable, since they are the true values from the given state until the end of the episode, we approximate them using the standard Bellman equation for Q-values. Under this setup, we find that inequality 18 holds in approximately 79% of cases across the 57 Atari games. These empirical findings indicate that although ˜4.1 is not theoretically guaranteed for $L_{2}$, it holds in the vast majority of the optimization steps in practice, which may explain the strong empirical performance of our method using the $L_{2}$ loss.
D.5 Empirical Evaluation of the $\delta$ Term Magnitude
In Theorem˜4.10 the $\delta$ term is ignored during optimization since it is independent of $Q$. To assess whether this omission introduces significant looseness in the causal bound, we conduct an empirical analysis on the Atari 57 benchmark using a DQN agent trained for 400K steps. Specifically, we plotted the magnitudes of the standard DQN loss, the $\delta$ term, and the SUFT OPE term throughout training. As discussed in Section˜D.4, the $\delta$ term involves the true values of $y_{0}$ and $y_{1}$, which are not directly observable. Thus, we approximate them using the standard Bellman equation for Q-values. The results, shown in Figure˜5, indicate that across all of the environments, the $\delta$ term remains consistently an order of magnitude lower than the standard DQN loss, and on the same order of magnitude as the SUFT OPE term. These findings support the decision to ignore the $\delta$ term during optimization, as it does not meaningfully loosen the causal bound.
Figure 5: Magnitude of the $\delta$ term, SUFT OPE term, and the standard DQN loss across the Atari 57 benchmark.
D.6 Atari 2600 57 Games Benchmark
Table 2: Atari 2600 random and average human rewards. The scores are sourced from [79].
Game Random Human
Alien 227.8 6371.3
Amidar 5.8 1540.4
Assault 222.4 628.9
Asterix 164.5 7536
Asteroids 871.3 36517.3
Atlantis 13463 26575
Bank Heist 14.2 644.5
Battle Zone 3560 33030
Beam Rider 254.6 14961
Berzerk 196.1 2237.5
Bowling 35.2 146.5
Boxing -1.5 9.6
Breakout 1.6 27.9
Centipede 2090.9 10321.9
Chopper Command 811 7387.8
Crazy Climber 10780.5 32667
Defender 2874.5 14296
Demon Attack 208.3 1971
Double Dunk -18.6 -16.4
Enduro 0 740.2
Fishing Derby -77.1 -38.7
Freeway 0.1 25.6
Frostbite 65.2 4202.8
Gopher 257.6 2311
Gravitar 173 3116
H.E.R.O. 1580.3 25839.4
Ice Hockey -11.2 0.5
James Bond 33.5 302.8
Kangaroo 100 2739
Krull 1598 2109.1
Kung-Fu Master 258.5 20786.8
Montezuma’s Revenge 25 4182
Ms. Pacman 307.3 6951.6
Name This Game 1747.8 6796
Phoenix 761.4 6686.2
Pit Fall -348.8 5998.9
Pong -20.7 14.6
Private Eye 662.8 64169.1
Q*Bert 163.9 12085
River Raid 1338.5 14382.2
Road Runner 200 6878
Robotank 2.2 8.9
Seaquest 215.5 40425.8
Skiing -15287.4 -4336.9
Solaris 2047.2 11032.6
Space Invaders 182.6 1464.9
Star Gunner 697 9528
Surround -10 5.4
Tennis -23.8 -8.3
Time Pilot 3568 5229.2
Tutankham 12.7 138.3
Up N Down 707.2 9896.1
Venture 18 1039
Video Pinball 20452 15641.1
Wizard of Wor 804 4556
Yars Revenge 3092.9 47135.2
Zaxxon 32.5 8443
Atari PPO Results (Using $L_{1}$} Loss):
Table 3: Training rewards comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 40 games, with statistically significant gains in 31 games. Both agents are using $L_{1}$} loss.
Game Baseline PPO SUFT PPO p-value
Alien 860.0 780.0 N/A
Amidar 184.0 133.0 N/A
Assault 631.0 692.0 9.15e-02
Asterix 346.0 568.0 1.17e-08
Asteroids 483.0 644.0 3.19e-07
Atlantis 36100.0 51300.0 4.50e-06
Bank Heist 37.5 51.5 6.14e-02
Battle Zone 8380.0 8880.0 6.07e-01
Beam Rider 450.0 526.0 2.34e-05
Berzerk 661.0 676.0 4.70e-01
Bowling 41.7 33.8 N/A
Boxing -10.0 15.1 1.35e-13
Breakout 14.4 10.4 N/A
Centipede 2860.0 3390.0 2.23e-06
Chopper Command 1350.0 1990.0 2.84e-05
Crazy Climber 26700.0 37800.0 4.67e-11
Defender 2850.0 4280.0 8.90e-08
Demon Attack 295.0 306.0 5.37e-01
Double Dunk -21.7 -18.1 2.19e-03
Enduro 13.1 67.8 3.18e-06
Fishing Derby -82.9 -79.5 2.95e-01
Freeway 20.2 24.6 6.60e-06
Frostbite 1610.0 261.0 N/A
Gopher 381.0 494.0 2.38e-05
Gravitar 318.0 211.0 N/A
H.E.R.O. 9810.0 8840.0 N/A
Ice Hockey -9.05 -5.86 2.65e-05
James Bond 324.0 466.0 3.54e-08
Kangaroo 3000.0 1080.0 N/A
Krull 4770.0 4360.0 N/A
Kung-Fu Master 1990.0 2130.0 1.05e-01
Montezuma’s Revenge 0.0 0.0 N/A
Ms. Pacman 1210.0 1770.0 1.55e-06
Name This Game 2480.0 4520.0 2.10e-13
Phoenix 734.0 1700.0 2.19e-11
Pit Fall -1.28 -4.73 N/A
Pong -16.5 -12.0 4.95e-03
Private Eye 84.3 25.7 N/A
Q*Bert 2080.0 689.0 N/A
River Raid 2560.0 2590.0 5.53e-01
Road Runner 2330.0 8600.0 4.66e-03
Robotank 4.88 4.1 N/A
Seaquest 576.0 689.0 4.14e-03
Skiing -30000.0 -18000.0 2.19e-04
Solaris 1850.0 1880.0 6.68e-01
Space Invaders 292.0 406.0 7.12e-06
Star Gunner 1030.0 1090.0 7.81e-04
Surround -9.7 -9.04 1.59e-02
Tennis -13.3 -15.6 N/A
Time Pilot 2790.0 3610.0 4.88e-05
Tutankham 82.1 102.0 2.85e-04
Up N Down 3090.0 19100.0 7.83e-06
Venture 16.0 8.0 N/A
Video Pinball 9030.0 14600.0 9.20e-06
Wizard of Wor 508.0 1060.0 2.09e-09
Yars Revenge 5230.0 10700.0 2.87e-12
Zaxxon 1040.0 271.0 N/A
Mean (%) 100.0 164.7 N/A
Figure 6: Learning curves comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{1}$} loss.
Figure 7: Log-scaled reward improvements comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 41 valid Atari games. SUFT outperforms the baseline agent in 28 out of the 41 valid games, achieving a reward improvement of over 100% in 24.4% of them and over 10% in 61% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{1}$} loss.
Atari PPO Results (Using $L_{2}$} Loss):
Table 4: Training rewards comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 50 games, with statistically significant gains in 40 games. Both agents are using $L_{2}$} loss.
Game Baseline PPO SUFT PPO p-value
Alien 860.0 861.0 3.48e-01
Amidar 184.0 184.0 N/A
Assault 631.0 1150.0 3.24e-07
Asterix 346.0 651.0 7.43e-08
Asteroids 483.0 586.0 9.18e-05
Atlantis 36100.0 59100.0 2.57e-09
Bank Heist 37.5 38.8 3.47e-01
Battle Zone 8380.0 11000.0 5.02e-05
Beam Rider 450.0 497.0 7.57e-03
Berzerk 661.0 782.0 3.34e-05
Bowling 41.7 43.2 3.73e-01
Boxing -10.0 1.75 2.10e-09
Breakout 14.4 17.3 1.90e-04
Centipede 2860.0 4120.0 3.28e-09
Chopper Command 1350.0 2940.0 4.37e-07
Crazy Climber 26700.0 40200.0 1.99e-10
Defender 2850.0 4070.0 3.08e-07
Demon Attack 295.0 634.0 2.31e-06
Double Dunk -21.7 -21.5 8.58e-01
Enduro 13.1 70.1 1.30e-10
Fishing Derby -82.9 -73.9 4.32e-06
Freeway 20.2 21.2 1.74e-02
Frostbite 1610.0 2750.0 5.59e-02
Gopher 381.0 547.0 1.71e-06
Gravitar 318.0 370.0 3.63e-01
H.E.R.O. 9810.0 11700.0 1.32e-01
Ice Hockey -9.05 -7.45 4.09e-04
James Bond 324.0 380.0 6.50e-05
Kangaroo 3000.0 5830.0 1.04e-01
Krull 4770.0 8980.0 1.36e-03
Kung-Fu Master 1990.0 2610.0 3.76e-03
Montezuma’s Revenge 0.0 0.0 N/A
Ms. Pacman 1210.0 1710.0 3.61e-03
Name This Game 2480.0 4410.0 7.45e-16
Phoenix 734.0 1720.0 1.05e-13
Pit Fall -1.28 -1.23 2.23e-01
Pong -16.5 -15.0 6.31e-03
Private Eye 84.3 76.0 N/A
Q*Bert 2080.0 5100.0 1.72e-03
River Raid 2560.0 3540.0 2.84e-11
Road Runner 2330.0 6320.0 1.51e-06
Robotank 4.88 4.29 N/A
Seaquest 576.0 1000.0 1.19e-07
Skiing -30000.0 -30000.0 N/A
Solaris 1850.0 2200.0 3.81e-01
Space Invaders 292.0 387.0 1.31e-04
Star Gunner 1030.0 1130.0 1.60e-08
Surround -9.7 -8.65 7.33e-04
Tennis -13.3 -13.6 N/A
Time Pilot 2790.0 3190.0 1.24e-02
Tutankham 82.1 129.0 1.79e-02
Up N Down 3090.0 6020.0 3.83e-03
Venture 16.0 5.0 N/A
Video Pinball 9030.0 11500.0 3.22e-03
Wizard of Wor 508.0 734.0 6.22e-04
Yars Revenge 5230.0 10100.0 1.36e-04
Zaxxon 1040.0 2310.0 7.88e-03
Mean (%) 100.0 197.34 N/A
Figure 8: Learning curves comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{2}$} loss.
Figure 9: Log-scaled reward improvements comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 42 valid Atari games. SUFT outperforms the baseline agent in 39 out of the 42 valid games, achieving a reward improvement of over 100% in 38.1% of them and over 10% in 83.3% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{2}$} loss.
Table 5: Training rewards comparison between the PPO agent using the additional SUFT OPE term with different $\lambda_{TF}$ values and the baseline agent without it across 57 Atari games. All agents are using $L_{2}$} loss.
Agent $\lambda_{TF}$ Mean (%)
PPO 0.0 100.0%
SUFT PPO 0.1 105.0%
0.2 116.05%
0.3 124.0%
0.4 126.99%
0.5 128.79%
0.6 137.06%
0.7 143.12%
0.8 146.56%
0.9 145.4%
1.0 151.72%
1.5 170.4%
2.0 167.33%
2.5 179.01%
3.0 184.66%
5.0 197.34%
10.0 198.28%
Atari DQN Results:
Table 6: Training rewards comparison between DQN agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 52 games, with statistically significant gains in 49 games. Both agents are using $L_{2}$} loss.
Game Baseline DQN, buffer 4K SUFT DQN, buffer 4K p-value
Alien 518.0 717.0 3.58e-07
Amidar 121.0 198.0 5.34e-06
Assault 1280.0 2250.0 4.10e-10
Asterix 872.0 1130.0 2.82e-03
Asteroids 294.0 450.0 2.31e-05
Atlantis 14000.0 19300.0 9.56e-05
Bank Heist 12.8 34.4 1.04e-05
Battle Zone 3350.0 4630.0 2.97e-04
Beam Rider 514.0 1050.0 6.13e-11
Berzerk 507.0 622.0 3.44e-06
Bowling 21.9 31.1 4.56e-02
Boxing -29.9 17.4 3.70e-15
Breakout 12.8 18.3 5.80e-03
Centipede 3160.0 3220.0 8.00e-01
Chopper Command 1150.0 1380.0 1.91e-03
Crazy Climber 21700.0 36000.0 3.05e-11
Defender 2510.0 2530.0 3.16e-01
Demon Attack 2240.0 4430.0 9.42e-11
Double Dunk -23.8 -22.9 1.96e-06
Enduro 48.2 169.0 3.24e-16
Fishing Derby -88.9 -39.8 2.55e-08
Freeway 19.0 21.2 1.39e-03
Frostbite 128.0 607.0 1.03e-03
Gopher 288.0 614.0 1.72e-06
Gravitar 136.0 284.0 2.10e-06
H.E.R.O. 2010.0 4110.0 1.56e-05
Ice Hockey -13.0 -6.62 1.19e-07
James Bond 34.0 363.0 1.24e-10
Kangaroo 460.0 5810.0 5.23e-07
Krull 315.0 3670.0 1.96e-14
Kung-Fu Master 6.0 2920.0 2.42e-05
Montezuma’s Revenge 0.0 0.0 N/A
Ms. Pacman 1060.0 1270.0 6.08e-05
Name This Game 2540.0 4250.0 6.56e-11
Phoenix 608.0 3120.0 3.82e-16
Pit Fall -163.0 -78.1 2.47e-03
Pong -20.2 -17.9 2.01e-07
Private Eye 133.0 154.0 7.51e-01
Q*Bert 908.0 2320.0 5.64e-06
River Raid 2410.0 3160.0 2.25e-08
Road Runner 991.0 10400.0 1.59e-10
Robotank 4.9 8.71 2.55e-07
Seaquest 393.0 911.0 2.56e-08
Skiing -30800.0 -30200.0 8.17e-04
Solaris 565.0 696.0 2.62e-04
Space Invaders 317.0 371.0 2.53e-04
Star Gunner 519.0 953.0 7.92e-11
Surround -9.11 -9.17 N/A
Tennis -18.9 -19.6 N/A
Time Pilot 1210.0 3350.0 4.85e-14
Tutankham 11.2 71.4 7.98e-12
Up N Down 3180.0 5340.0 1.27e-03
Venture 4.0 0.0 N/A
Video Pinball 9360.0 8890.0 N/A
Wizard of Wor 235.0 421.0 4.94e-09
Yars Revenge 3060.0 8280.0 7.44e-08
Zaxxon 0.0 2710.0 1.23e-07
Mean (%) 100.0 383.19 N/A
Figure 10: Learning curves comparison between DQN agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{2}$} loss.
Figure 11: Log-scaled reward improvements comparison between DQN agent using the additional SUFT OPE term and the baseline agent without it across 29 valid Atari games. SUFT outperforms the baseline agent in 28 out of the 29 valid games, achieving a reward improvement of over 100% in 48.3% of them and over 10% in 93.1% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{2}$} loss.
Table 7: Training rewards comparison between DQN agent using the additional SUFT OPE term with a small 4K buffer and the baseline agent without the SUFT OPE term with a larger 100K buffer across 57 Atari games. SUFT surpasses the baseline agent in 52 games, with statistically significant gains in 47 games. Both agents are using $L_{2}$} loss.
Game Baseline DQN, buffer 100K SUFT DQN, buffer 4K p-value
Alien 438.0 717.0 3.14e-09
Amidar 118.0 198.0 2.20e-07
Assault 1120.0 2250.0 6.54e-10
Asterix 914.0 1130.0 4.06e-03
Asteroids 275.0 450.0 4.69e-06
Atlantis 7970.0 19300.0 7.59e-09
Bank Heist 9.9 34.4 1.01e-11
Battle Zone 3100.0 4630.0 9.71e-04
Beam Rider 561.0 1050.0 7.44e-09
Berzerk 458.0 622.0 1.55e-05
Bowling 21.8 31.1 4.58e-02
Boxing -28.7 17.4 2.25e-15
Breakout 17.7 18.3 6.94e-01
Centipede 3130.0 3220.0 7.99e-01
Chopper Command 648.0 1380.0 1.30e-09
Crazy Climber 30400.0 36000.0 1.38e-04
Defender 2440.0 2530.0 4.85e-02
Demon Attack 3220.0 4430.0 8.29e-08
Double Dunk -23.9 -22.9 8.65e-07
Enduro 66.8 169.0 5.22e-16
Fishing Derby -84.1 -39.8 4.29e-07
Freeway 18.9 21.2 2.32e-03
Frostbite 159.0 607.0 1.95e-03
Gopher 119.0 614.0 3.64e-09
Gravitar 107.0 284.0 4.17e-09
H.E.R.O. 2320.0 4110.0 8.37e-05
Ice Hockey -13.2 -6.62 2.20e-08
James Bond 43.5 363.0 1.41e-10
Kangaroo 534.0 5810.0 5.25e-07
Krull 336.0 3670.0 2.71e-14
Kung-Fu Master 13.0 2920.0 2.17e-05
Montezuma’s Revenge 0.0 0.0 N/A
Ms. Pacman 1060.0 1270.0 3.69e-03
Name This Game 2990.0 4250.0 3.66e-10
Phoenix 732.0 3120.0 1.12e-15
Pit Fall -114.0 -78.1 5.41e-01
Pong -20.6 -17.9 7.89e-09
Private Eye 110.0 154.0 6.69e-01
Q*Bert 1130.0 2320.0 3.81e-05
River Raid 2200.0 3160.0 3.60e-10
Road Runner 909.0 10400.0 1.04e-10
Robotank 4.44 8.71 4.26e-05
Seaquest 391.0 911.0 1.08e-08
Skiing -30700.0 -30200.0 8.64e-03
Solaris 515.0 696.0 2.80e-05
Space Invaders 366.0 371.0 3.07e-01
Star Gunner 351.0 953.0 1.21e-12
Surround -8.69 -9.17 N/A
Tennis -17.5 -19.6 N/A
Time Pilot 1100.0 3350.0 1.47e-14
Tutankham 9.4 71.4 4.66e-18
Up N Down 2160.0 5340.0 6.09e-05
Venture 4.0 0.0 N/A
Video Pinball 9890.0 8890.0 N/A
Wizard of Wor 292.0 421.0 2.39e-07
Yars Revenge 3530.0 8280.0 9.75e-08
Zaxxon 0.0 2710.0 1.19e-07
Mean (%) 100.0 437.05 N/A
Table 8: Training rewards comparison between DQN agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 49 games, with statistically significant gains in 44 games. Both agents are using a 100k buffer and $L_{2}$} loss.
Game Baseline DQN, buffer 100K SUFT DQN, buffer 100K p-value
Alien 438.0 686.0 1.45e-08
Amidar 118.0 235.0 6.49e-11
Assault 1120.0 2170.0 3.18e-10
Asterix 914.0 1240.0 3.19e-05
Asteroids 275.0 430.0 1.03e-05
Atlantis 7970.0 12400.0 5.88e-05
Bank Heist 9.9 27.7 8.87e-04
Battle Zone 3100.0 4270.0 5.45e-05
Beam Rider 561.0 1010.0 1.61e-09
Berzerk 458.0 625.0 2.38e-05
Bowling 21.8 21.5 N/A
Boxing -28.7 -1.34 1.10e-10
Breakout 17.7 18.1 5.37e-01
Centipede 3130.0 4820.0 1.47e-02
Chopper Command 648.0 957.0 9.89e-04
Crazy Climber 30400.0 35500.0 6.89e-05
Defender 2440.0 2590.0 3.53e-02
Demon Attack 3220.0 5220.0 1.85e-10
Double Dunk -23.9 -22.5 1.34e-11
Enduro 66.8 153.0 2.45e-15
Fishing Derby -84.1 -43.9 1.87e-08
Freeway 18.9 21.1 4.23e-03
Frostbite 159.0 33.3 N/A
Gopher 119.0 666.0 9.22e-07
Gravitar 107.0 286.0 2.30e-04
H.E.R.O. 2320.0 431.0 N/A
Ice Hockey -13.2 -5.29 6.87e-10
James Bond 43.5 372.0 2.39e-06
Kangaroo 534.0 5670.0 4.11e-14
Krull 336.0 1780.0 2.90e-05
Kung-Fu Master 13.0 3190.0 5.18e-09
Montezuma’s Revenge 0.0 0.0 N/A
Ms. Pacman 1060.0 1350.0 9.98e-04
Name This Game 2990.0 3680.0 9.95e-06
Phoenix 732.0 2700.0 4.01e-15
Pit Fall -114.0 -355.0 N/A
Pong -20.6 -18.3 4.76e-06
Private Eye 110.0 82.4 N/A
Q*Bert 1130.0 2850.0 1.01e-05
River Raid 2200.0 2950.0 4.95e-11
Road Runner 909.0 9240.0 1.54e-13
Robotank 4.44 8.95 3.82e-05
Seaquest 391.0 574.0 5.20e-04
Skiing -30700.0 -30200.0 1.39e-02
Solaris 515.0 653.0 1.07e-03
Space Invaders 366.0 421.0 4.64e-06
Star Gunner 351.0 387.0 3.01e-01
Surround -8.69 -8.55 4.83e-01
Tennis -17.5 -18.5 N/A
Time Pilot 1100.0 3570.0 1.43e-15
Tutankham 9.4 15.0 2.77e-01
Up N Down 2160.0 3020.0 2.25e-03
Venture 4.0 10.0 1.46e-02
Video Pinball 9890.0 8970.0 N/A
Wizard of Wor 292.0 423.0 4.12e-03
Yars Revenge 3530.0 6900.0 7.48e-02
Zaxxon 0.0 1020.0 1.08e-02
Mean (%) 100.0 413.64 N/A
Atari Vanilla DQN Results:
Table 9: Training rewards comparison between Vanilla DQN agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 49 games, with statistically significant gains in 33 games. Both agents are using $L_{2}$} loss.
Game Baseline Vanilla DQN, buffer 4K SUFT Vanilla DQN, buffer 4K p-value
Alien 657.0 736.0 1.37e-02
Amidar 163.0 201.0 1.95e-04
Assault 904.0 1600.0 7.33e-11
Asterix 680.0 1030.0 8.30e-09
Asteroids 408.0 423.0 5.36e-01
Atlantis 10300.0 19000.0 6.54e-08
Bank Heist 29.4 35.2 8.01e-02
Battle Zone 11600.0 12700.0 1.10e-01
Beam Rider 909.0 1440.0 8.25e-08
Berzerk 613.0 653.0 2.35e-03
Bowling 26.5 25.1 N/A
Boxing 0.31 17.8 1.10e-10
Breakout 16.0 18.0 5.85e-01
Centipede 3430.0 3680.0 2.13e-02
Chopper Command 1390.0 1700.0 1.84e-04
Crazy Climber 38700.0 43500.0 1.76e-05
Defender 3290.0 4000.0 5.24e-04
Demon Attack 1490.0 2790.0 6.15e-12
Double Dunk -22.9 -23.0 N/A
Enduro 156.0 207.0 3.40e-08
Fishing Derby -76.0 -60.8 4.17e-05
Freeway 24.9 25.2 9.93e-01
Frostbite 272.0 348.0 1.32e-01
Gopher 362.0 654.0 8.02e-06
Gravitar 221.0 224.0 5.56e-01
H.E.R.O. 4590.0 8700.0 1.55e-02
Ice Hockey -10.8 -9.41 2.94e-02
James Bond 402.0 428.0 1.80e-01
Kangaroo 1780.0 4360.0 1.95e-02
Krull 3650.0 4440.0 8.03e-05
Kung-Fu Master 2880.0 2660.0 N/A
Montezuma’s Revenge 0.0 0.0 N/A
Ms. Pacman 1560.0 1520.0 N/A
Name This Game 3060.0 4170.0 2.79e-09
Phoenix 2110.0 2720.0 7.66e-07
Pit Fall -35.2 -42.2 N/A
Pong -15.6 -16.9 N/A
Private Eye 66.4 72.6 5.77e-01
Q*Bert 1510.0 2100.0 1.69e-03
River Raid 2970.0 3420.0 4.87e-06
Road Runner 16900.0 17700.0 5.00e-01
Robotank 5.62 7.02 1.02e-04
Seaquest 751.0 1120.0 1.08e-07
Skiing -30300.0 -30200.0 7.50e-02
Solaris 915.0 683.0 N/A
Space Invaders 374.0 424.0 5.24e-04
Star Gunner 829.0 978.0 1.53e-05
Surround -9.42 -9.33 5.56e-01
Tennis -19.5 -18.0 2.67e-02
Time Pilot 2880.0 2910.0 5.70e-01
Tutankham 53.5 106.0 1.03e-07
Up N Down 4640.0 6970.0 8.16e-05
Venture 9.0 18.0 1.29e-01
Video Pinball 13700.0 14200.0 2.84e-01
Wizard of Wor 365.0 424.0 1.79e-03
Yars Revenge 7260.0 10100.0 1.42e-07
Zaxxon 112.0 932.0 1.96e-01
Mean (%) 100.0 228.12 N/A
Figure 12: Learning curves comparison between Vanilla DQN agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{2}$} loss.
Figure 13: Log-scaled reward improvements comparison between Vanilla DQN agent using the additional SUFT OPE term and the baseline agent without it across 43 valid Atari games. SUFT outperforms the baseline agent in 39 out of the 43 valid games, achieving a reward improvement of over 100% in 25.6% of them and over 10% in 79.1% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{2}$} loss.
Atari Double DQN Results:
Table 10: Training rewards comparison between the Double DQN agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 48 games, with statistically significant gains in 37 games. Both agents are using $L_{2}$} loss.
Game Baseline DDQN, buffer 4K SUFT DDQN, buffer 4K p-value
Alien 773.0 806.0 7.88e-01
Amidar 152.0 177.0 8.27e-03
Assault 781.0 1440.0 3.84e-12
Asterix 676.0 939.0 1.03e-08
Asteroids 357.0 365.0 1.30e-01
Atlantis 13900.0 18100.0 1.14e-03
Bank Heist 29.2 36.1 1.30e-03
Battle Zone 10900.0 14000.0 1.16e-04
Beam Rider 719.0 1230.0 1.03e-12
Berzerk 567.0 668.0 5.01e-08
Bowling 27.8 27.0 N/A
Boxing -1.73 18.4 8.10e-11
Breakout 19.3 16.3 N/A
Centipede 3390.0 3430.0 6.14e-01
Chopper Command 1460.0 1870.0 4.47e-05
Crazy Climber 41200.0 42200.0 3.93e-01
Defender 2980.0 4140.0 8.61e-10
Demon Attack 369.0 2240.0 3.09e-15
Double Dunk -23.6 -23.4 2.41e-02
Enduro 143.0 195.0 3.43e-13
Fishing Derby -81.4 -67.8 4.39e-07
Freeway 25.2 21.3 N/A
Frostbite 344.0 265.0 N/A
Gopher 595.0 658.0 4.17e-01
Gravitar 100.0 134.0 2.00e-02
H.E.R.O. 4020.0 8870.0 2.06e-03
Ice Hockey -7.7 -8.15 N/A
James Bond 316.0 416.0 7.84e-04
Kangaroo 2430.0 3280.0 1.94e-01
Krull 3830.0 4400.0 9.09e-05
Kung-Fu Master 2470.0 3680.0 3.30e-02
Montezuma’s Revenge 0.0 0.0 N/A
Ms. Pacman 1560.0 1650.0 3.26e-02
Name This Game 3140.0 4140.0 7.71e-08
Phoenix 1080.0 2220.0 1.05e-10
Pit Fall -14.2 -27.5 N/A
Pong -15.9 -15.8 6.54e-01
Private Eye -45.2 49.4 8.98e-01
Q*Bert 1260.0 1960.0 6.31e-05
River Raid 2960.0 3420.0 1.56e-07
Road Runner 15700.0 17000.0 4.29e-01
Robotank 6.43 7.47 6.65e-03
Seaquest 474.0 938.0 9.91e-12
Skiing -30300.0 -30200.0 2.87e-01
Solaris 1620.0 599.0 N/A
Space Invaders 377.0 446.0 3.52e-05
Star Gunner 970.0 1100.0 5.04e-06
Surround -9.62 -9.64 N/A
Tennis -23.6 -20.3 7.08e-05
Time Pilot 2130.0 2640.0 3.75e-05
Tutankham 64.1 107.0 1.56e-05
Up N Down 3940.0 6950.0 1.23e-04
Venture 6.0 18.0 4.54e-02
Video Pinball 13400.0 15400.0 1.23e-02
Wizard of Wor 349.0 447.0 3.76e-05
Yars Revenge 8610.0 10400.0 2.31e-06
Zaxxon 4.0 10.0 1.71e-01
Mean (%) 100.0 226.91 N/A
Figure 14: Learning curves comparison between the Double DQN agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{2}$} loss.
Figure 15: Log-scaled reward improvements comparison between the Double DQN agent using the additional SUFT OPE term and the baseline agent without it across 40 valid Atari games. SUFT outperforms the baseline agent in 35 out of the 40 valid games, achieving a reward improvement of over 100% in 20% of them and over 10% in 72.5% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{2}$} loss.
Table 11: Training rewards comparison between our causal bound optimization method and an inertia regularizer. We compare a Double DQN agent with the additional SUFT OPE term to one using an inertia regularizer across 57 Atari games. While SUFT significantly improves the baseline agent’s performance with 226.91% mean reward ratio, the inertia regularizer harms the baseline agent’s performance with only 83.21% mean reward ratio. All agents are using $L_{2}$} loss and a 4K buffer size.
Game Baseline DDQN Inertia Regularizer DDQN SUFT DDQN
Alien 773.0 762.0 806.0
Amidar 152.0 144.0 177.0
Assault 781.0 670.0 1440.0
Asterix 676.0 666.0 939.0
Asteroids 357.0 278.0 365.0
Atlantis 13900.0 3970.0 18100.0
Bank Heist 29.2 21.8 36.1
Battle Zone 10900.0 11200.0 14000.0
Beam Rider 719.0 707.0 1230.0
Berzerk 567.0 628.0 668.0
Bowling 27.8 28.8 27.0
Boxing -1.73 -3.91 18.4
Breakout 19.3 8.85 16.3
Centipede 3390.0 3840.0 3430.0
Chopper Command 1460.0 1120.0 1870.0
Crazy Climber 41200.0 42800.0 42200.0
Defender 2980.0 2650.0 4140.0
Demon Attack 369.0 191.0 2240.0
Double Dunk -23.6 -23.6 -23.4
Enduro 143.0 107.0 195.0
Fishing Derby -81.4 -84.9 -67.8
Freeway 25.2 21.2 21.3
Frostbite 344.0 273.0 265.0
Gopher 595.0 376.0 658.0
Gravitar 100.0 138.0 134.0
H.E.R.O. 4020.0 5060.0 8870.0
Ice Hockey -7.7 -6.6 -8.15
James Bond 316.0 186.0 416.0
Kangaroo 2430.0 1110.0 3280.0
Krull 3830.0 2700.0 4400.0
Kung-Fu Master 2470.0 1520.0 3680.0
Montezuma’s Revenge 0.0 0.0 0.0
Ms. Pacman 1560.0 1510.0 1650.0
Name This Game 3140.0 2420.0 4140.0
Phoenix 1080.0 640.0 2220.0
Pit Fall -14.2 -25.5 -27.5
Pong -15.9 -17.3 -15.8
Private Eye -45.2 -11.9 49.4
Q*Bert 1260.0 1180.0 1960.0
River Raid 2960.0 2550.0 3420.0
Road Runner 15700.0 11000.0 17000.0
Robotank 6.43 6.51 7.47
Seaquest 474.0 306.0 938.0
Skiing -30300.0 -26500.0 -30200.0
Solaris 1620.0 1860.0 599.0
Space Invaders 377.0 400.0 446.0
Star Gunner 970.0 980.0 1100.0
Surround -9.62 -9.37 -9.64
Tennis -23.6 -15.1 -20.3
Time Pilot 2130.0 1920.0 2640.0
Tutankham 64.1 75.8 107.0
Up N Down 3940.0 2930.0 6950.0
Venture 6.0 4.0 18.0
Video Pinball 13400.0 12400.0 15400.0
Wizard of Wor 349.0 215.0 447.0
Yars Revenge 8610.0 7990.0 10400.0
Zaxxon 4.0 12.0 10.0
Mean (%) 100.0 83.21 226.91
D.7 MuJoCo Five Environments Benchmark
MuJoCo PPO Results:
Table 12: Training rewards comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 5 MuJoCo environments. SUFT surpasses the baseline agent in 4 out of 5 environments, with statistically significant gains in 2 environments. Both agents are using $L_{2}$} loss.
Game Baseline PPO SUFT PPO p-value
Ant 1110.0 1570.0 1.65e-02
Half Cheetah 1520.0 1540.0 7.28e-01
Hopper 1880.0 2120.0 3.16e-03
Humanoid 464.0 464.0 N/A
Walker2d 1850.0 1860.0 8.33e-01
Mean (%) 100.0 111.21 N/A
Figure 16: Learning curves comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 5 MuJoCo environments. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{2}$} loss.
MuJoCo SAC Results:
Table 13: Training rewards comparison between the SAC agent using the additional SUFT OPE term and the baseline agent without it across 5 MuJoCo environments. SUFT surpasses the baseline agent in 4 out of 5 environments, with statistically significant gains in 3 environments. Both agents are using $L_{2}$} loss.
Game Baseline SAC, buffer 4K SUFT SAC, buffer 4K p-value
Ant 1250.0 1020.0 N/A
Half Cheetah 4440.0 6290.0 3.94e-01
Hopper 978.0 1910.0 1.46e-06
Humanoid 638.0 1100.0 2.16e-02
Walker2d 316.0 1680.0 8.16e-11
Mean (%) 100.0 224.53 N/A
Figure 17: Learning curves comparison between the SAC agent using the additional SUFT OPE term and the baseline agent without it across 5 MuJoCo environments. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{2}$} loss.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract and introduction, we highlight the main problem, our method, and the empirical results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Section˜4, we describe the assumptions of our work, and in particular ˜4.1, which does not apply to the $L_{2}$ loss. In our experiments (Section˜5), we evaluate our method on $L_{2}$ loss.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In Section˜4, we detailed the relevant assumptions for all theoretical results. We provide all proofs in Appendix˜B.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We use open-source environments and an RL library as detailed in the Implementation Details paragraph in Section˜5. Furthermore, we fully detailed all the experimental information in the Training Details paragraph in Section˜5. Throughout Section˜4 and Appendix˜B, we thoroughly detail how to implement our method. We also provide all the evaluation formulas in Appendix˜D. In addition, we also provide all of the code as supplementary material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the code as supplementary material and fully detail the implementation and training details in Section˜5. We also provide all the evaluation formulas in Appendix˜D.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We fully detailed all the experimental information in the Training Details and Implementation Details paragraphs in Section˜5.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All experiments are independently trained for 10 different random seeds per agent-environment configuration, and we take the median result. We report mean performances for all of the environments per domain. We also report environment-specific statistical significance results by conducting a $t$-test.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We give details on the computational time and resources in Appendix˜D.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper proposes a method to improve sample efficiency in reinforcement learning. As foundational research, this work has no direct paths to negative societal impacts besides those that can arise from improved reinforcement learning agents.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release high-risk data or models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all creators whose code we used in our experiments in Section˜5.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the code as supplementary material and include a README file that explains how to run it.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This paper uses LLMs only for editing purposes and conforms to the NeurIPS LLM policy.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.